### v1

# Dispersion on Time-Varying Graphs

Abstract
The dispersion problem involves the coordination of $k\leq n$ agents on a graph of size $n$ to reach a configuration where at each node at most one agent can be present. It is a well-studied problem on static graphs. Also, this problem is studied on dynamic graphs with $n$ nodes where at each discrete time step the graph is a connected sub-graph of the complete graph $K_{n}$. An optimal algorithm is provided assuming global communication and 1-hop visibility of the agents. How this problem pans out on Time-Varying Graphs (TVG) is kept as an open question in the literature. In this work we study this problem on TVG where at each discrete time step the graph is a connected sub-graph of an underlying graph $G$ (known as a footprint) consisting of $n$ nodes. We have the following results even if only at most one edge from $G$ is missing in the connected sub-graph at any time step, and all agents start from a rooted initial configuration. Even with unlimited memory at each agent and 1-hop visibility, it is impossible to solve dispersion for $n$ co-located agents on a TVG in the local communication model. Furthermore, even with unlimited memory at each agent but without 1-hop visibility, it is impossible to achieve dispersion for $n$ co-located agents in the global communication model. From the positive side, the existing algorithm for dispersion on dynamic graphs with the assumptions of global communication and 1-hop visibility works on TVGs as well. This fact and the impossibility results push us to come up with a modified definition of the dispersion problem on TVGs, as one needs to start with more than $n$ agents if the objective is to drop the strong assumptions of global communication and 1-hop visibility. Then, we provide an algorithm to solve the modified dispersion problem on TVG starting with $n+1$ agents with $O(\log n)$ memory per agent while dropping both the assumptions of global communication and 1-hop visibility. This algorithm is optimal with respect to the number of agents (due to the impossibilities) and memory per agent.
Keywords:
Dispersion Mobile agents Distributed algorithms Time-Varying Graph Deterministic algorithms
1 Introduction
The dispersion problem, introduced in [2] on static graphs, involves the coordination of $k\leq n$ agents on the graph of size $n$ to reach a configuration with the constraint that at most one agent can be present at any node. This problem is generally applicable to real-world scenarios in which agents must coordinate and share resources located at various places. The goal is to minimize the total cost of solving the problem while ensuring that the cost of agents moving around on the graph is much lower than the cost of having more than one agent share a resource. The dispersion problem is also connected to three other problems: (i) scattering or uniform deployment of agents on a graph [4], (ii) exploration of agents on a graph [5], and (iii) load balancing on a graph [3]. The scattering problem involves coordinating and moving $k\leq n$ agents on a given graph with $n$ nodes so that they are uniformly distributed over the graph. When $k=n$, the scattering problem is equivalent to the dispersion problem. The problem of collaborative exploration requires agents to visit all the nodes of a given graph while coordinating with one another. This problem is related to dispersion as if $n$ agents need to disperse, they must explore the graph. Any solution to the dispersion problem can be applied to collaborative exploration as long as the assumptions and model parameters remain the same. Load balancing on graphs is a problem where resources need to be evenly distributed as much as possible among nodes on a graph in the least possible number of rounds. Typically, nodes control the allocation of resources, and there are constraints on how much load can pass through an edge in a round. This problem is similar to dispersion if we consider agents as resources.
There are several works on dispersion [2, 7, 8, 14, 11, 15], and most of them deal with static graphs (graphs where the topology of the graph does not change with time). The addition of dynamism to a network makes the dispersion problem more challenging as well as practical. Agents may have to complete their tasks in environments that are also changing. In this work, we aim to solve the dispersion problem on dynamic graphs. Below we discuss the existing works on dispersion.
1.1 Related work
The problem of dispersion has been extensively studied in the literature [2, 7, 8, 14, 11, 15]. Augustine et al. in [2] introduced the dispersion problem. For the case where $k=n$, in their work [2], the authors present a memory lower bound of $\Omega(\log n)$ bits for each agent and a time lower bound of $\Omega(D)$ ($\Omega(n)$ on arbitrary graphs) for any deterministic algorithm on a given graph, where $D$ represents the diameter of the graph. In [2], authors have given two algorithms on arbitrary graphs for $k=n$: (i) algorithm with $O(m\cdot n)$ time using $O(\log n)$ bits at each agent and (ii) algorithm with $O(m)$ time using $O(n\cdot\log n)$ bits memory at each agent, where $m$ is the number of edges in the graph. Kshemkalyani et al. [7] provide an $\Omega(k)$ time lower bound for dispersion on arbitrary graphs for any $k\leq n$. In [7], authors have given three algorithms for arbitrary graphs for $k\leq n$: (i) $O(m)$ time using $O(k\cdot\log\Delta)$ memory at each agent, (ii) $O(\Delta^{D})$ time using $O(D\cdot\log\Delta)$ memory at each agent, and (iii) $O(m\cdot k)$ time using $O(\log(max\{k,\,\Delta\}))$ memory at each agent, where $\Delta$ is the maximum degree of the graph. Recently, Kshemkalyani et al. [8] provided an $O(\min\{m,\,k\cdot\Delta\}\cdot\log k)$-time with $O(\log n)$ memory per agent algorithm for arbitrary graphs, provided that the parameters $m$, $k$, $\Delta$ are known to agents. Kshemkalyani et al. [11] provided an algorithm that solves dispersion in $O(min\{m,\,k\cdot\Delta\})$ time with $\Theta(\log(k+\Delta))$ bits at each agent improving the time bound of the best previously known algorithm by $O(\log l)$ where $l\leq\frac{k}{2}$ and matching asymptotically the single-source depth-first search(DFS) traversal bounds [8]. Shintaku et al. [14] further present a dispersion algorithm that does not require any global knowledge and solves the dispersion problem on arbitrary graphs in $O(\min\{m,\,k\cdot\Delta\}\cdot\log l)$ rounds using $\Theta(\log(k+\Delta))$ bits of memory at each agent. Recently, Sudo et al. [15] presented even faster algorithms for both rooted and arbitrary configurations. For the rooted setting, they provide two algorithms. The first algorithm achieves dispersion in $O(k\cdot\log min\{k,\,\Delta\})$ time using $O(\log\Delta)$ bits of memory per agent. The second algorithm is time optimal, i.e., $O(k)$ time using $O(\Delta)$ memory per agent. The algorithm provided for the general setting requires $O(k\cdot\log k\cdot\log min\{k,\,\Delta\})$ using $O(\log(k+\Delta))$ memory per agent.
Recently, researchers in the field of distributed computing started studying dynamic graphs, where the topological changes are not sporadic or anomalous, but rather inherent to the nature of the network. Such a dynamic graph model was developed by Kuhn et al. [12] in 2010. Dispersion on the dynamic graphs is studied by Agarwalla et al. [1] and Kshemkalyani et al. [10]. In [1], the authors consider the following dynamic graph model on ring $\mathcal{R}$. In [1], the adversary can delete at most on edge from $\mathcal{R}$ at each round. Based on this dynamic graph model, they develop several deterministic algorithms for agents to achieve dispersion on a ring. In [10], the authors use a weaker dynamic graph model where the adversary can add or remove edges keeping the dynamic graph connected i.e., at any round, the dynamic graph $\mathcal{G}$ is connected. The authors show that it is impossible to solve dispersion on the dynamic graph $\mathcal{G}$ in the local communication model, even if $1$-hop visibility is available to the agents and each agent has unlimited memory. Also, it is impossible to solve dispersion on the dynamic graph $\mathcal{G}$ in the global communication model without $1$-hop visibility, even with unlimited memory at each agent. They provided an asymptotically optimal $\Theta(k)$ rounds algorithm in the global communication model with $1$-hop visibility.
The authors [10] ask the following question on the behavior of the dispersion problem on time-varying graphs (TVG) as TVG exhibits a stronger graph model than their model. It is easy to observe that the optimal algorithm of [10] also works for TVGs (as the graph model of TVG is stronger) in the global communication model with $1$-hop visibility. Due to the same reason, their impossibility results do not hold here on TVGs. In this work, we study the dispersion problem on TVG and provide impossibility results on the dispersion of $k\leq n$ agents in TVG, even from a rooted initial configuration, in the absence of either $1$-hop visibility or global communication. We further provide an algorithm to solve dispersion with $n+1$ agents from a rooted initial configuration on TVG dropping their assumptions of $1$-hop visibility and global communication. Below we provide our model description and the problem definition.
1.2 Model and the problem
Time-varying graph model: The system is represented as a time-varying graph (TVG), denoted by $\mathcal{G}=(V,E,\mathbb{T},\rho)$, where, $V$ represents the set of nodes, $E$ represents a set of edges, $\mathbb{T}$ represents the temporal domain, which is defined to be $\mathbb{Z}^{+}$ as in this model we consider discrete time steps, and $\rho:E\times\mathbb{T}\rightarrow\{0,1\}$ tells whether a particular edge is available at a given time. The graph $G=(V,\,E)$ is the underlying static graph of the dynamic graph $\mathcal{G}$, also termed as a footprint of $\mathcal{G}$. We denote $\delta_{v}$ as the degree of node $v$ in the footprint $G$, and $\Delta$ as the maximum degree in the footprint $G$. The graph $G$ is undirected and unweighted. The nodes of $G$ are anonymous (i.e., they have no IDs). The graph $G$ is port-labelled, i.e., each incident edge of a node $v$ is assigned a locally unique port number in $\{0,1,2,3,\cdots,\delta_{v}-1\}$. Specifically, each undirected edge $(u,\,v)$ has two labels, denoted as its port numbers at $u$ and $v$. Port numbering is local, i.e., there is no relationship between the port numbers at $u$ and $v$. We call a node as hole in round $t$ if there is no agent at that node in round $t$. We call a node a multinode in round $t$ if there are at least two agents present at that node in round $t$. We consider there is an adversary that can remove at most one edge from the footprint $G$ keeping $\mathcal{G}$ connected at each round.
Agent: We consider $k$ agents to be present initially at nodes of the graph $G$. Each agent has a unique identifier assigned from the range $[1,\,k]$. Each agent knows its ID. Agents do not have any prior knowledge of $k$, or any other graph parameters like $n$ and $\Delta$. The agents are equipped with $O(\log k)$ memory. An agent residing at a node (say $v$) in round $t$, knows $\delta_{v}$ in the footprint $G$ and all associated ports corresponding to node $v$ in $G$; however, the agent does not understand if any incident edge of node $v$ is missing in $\mathcal{G}$ in round $t$. To be more precise, agent $a$ being at node $v$ in round $t$ and $e_{v}$ being any edge associated with $v$, $a$ does not know the value of $\rho(e_{v},\,t)$.
The algorithm runs in synchronous rounds. In each round $t$, an agent $a_{i}$ performs one Communicate-Compute-Move (CCM) cycle as follows:
  • Communicate: Agent $a_{i}$ at node $v_{i}$ can communicate with other agents $a_{j}$ present at the same node $v_{i}$ or present at any different node $v_{j}$, depending on the communication model used. The agent also understands whether it had a successful or an unsuccessful move in the last round.
  • Compute: Based on the information the agent has, the agent computes the port through which it will move or decides not to move at all.
  • Move: Agent moves via the computed port or stays at its current node.
Communication model: There are two communication models in a distributed setting: local and global. In the local communication model [2, 7], an agent located at a particular graph node can communicate with other agents present at the same node. In contrast, the global communication model [6, 10] allows an agent at a graph node to communicate with any other agent in the graph, even if they are located at different nodes. In nature, this type of communication happens through the links of the graph. We use the local communication model in our algorithm.
Visibility model: In this work, we use two types of visibility models: zero-hop visibility and $1$-hop visibility. In the zero-hop visibility model [2, 8, 9], an agent located at a particular node (say $v\in\mathcal{G}$) can see agents present at node $v$ in round $r$, and port numbers associated with node $v$. In $1$-hop visibility [1, 13], an agent (say $a_{i}$) located at a particular node (say $v\in\mathcal{G}$) can see any node (along with agent(s) present at that node, if any) that is the neighbours of $v$ in $\mathcal{G}$. If the adversary deletes an edge associated with a node $u$ with port $p$ at the start of a round $r$, the agent(s) present there cannot understand that an edge has been deleted in the zero-hop visibility model. However, if the agent tries to move through that edge in this round, it cannot make the move. In the 1-hop visibility model, if such an edge is deleted in a round, then the agents present there can see all the neighbours of $v$ that are connected to $v$ via the edges with the port number $\neq p$ at $v$ because according to the 1-hop visibility model, the agents can see the sub-graph induced by the nodes that are 1-hop away in this round. Therefore, agents can also recognize that no node is connected through port $p$, and therefore it can understand that the edge corresponding to port $p$ at $v$ is deleted at the start of this round.
Now we are ready to provide our problem definition.
Dispersion on TVG: Let $\mathcal{G}$ be a TVG. Let $k$ agents be at the nodes of the graph $G$. At each round, the adversary can remove at most one edge from the footprint $G$ keeping $\mathcal{G}$ connected. If $k\leq n$, the agents reposition themselves such that each agent is at a distinct node of $G$; else, if $k>n$, the agents reposition themselves such that each node of $G$ contains at least one agent.
1.3 Our contribution
In this paper, we provide the following results.
  • It is impossible to solve Dispersion on TVG with $k=n$ agents considering 1-hop visibility but without global communication, even if the agents start from a rooted configuration and have infinite memory (refer Section 2).
  • It is impossible to solve Dispersion on TVG with $k=n$ agents considering global communication but without 1-hop visibility, even if the agents start from a rooted configuration and have infinite memory (refer Section 2).
  • We present an algorithm (refer Section 3) that solves Dispersion on TVG with $n+1$ initially co-located agents in $O(n^{4}\cdot\log n)$ rounds that require $O(\log n)$ memory per agent considering local communication and zero visibility. This algorithm is optimal with respect to the number of agents given the impossibility results. Also, this algorithm is memory-optimal.
2 Impossibility Results
In this section, we show that it is impossible to solve dispersion deterministically in our model if either global communication or 1-hop visibility is omitted, even if agents start from a rooted configuration and have infinite memory. Recall that according to our model, the agents at a node $v$ cannot know whether an incident edge to node $v$ is missing or not.
Theorem 2.1
(Impossibility without global communication) It is impossible to solve Dispersion on TVG with $k=n$ agents for any $n\geq 6$ with the agents having $1$-hop visibility but no global communication even if agents are initially co-located and have infinite memory.
Proof
We prove the above theorem by contradiction. Let us consider a ring $\mathcal{R}$ of size $n$ and $n$ many agents be co-located at some node of the ring $\mathcal{R}$. Suppose an algorithm $\mathcal{A}$ solves Dispersion on TVG. If agents are not in the dispersed configuration at the beginning of round $r$, it implies that there is at least one multinode at the beginning of round $r$. Our idea is to show that there must exist a multinode in the beginning round $r+1$. Observe that, if there is some multinode of size at least 4 at the beginning of round $r$ then there must exist a multinode at the beginning of the next round. Based on the size of the multinode, we categorise this into two cases.
  • Case 1: At the beginning round $r$, there is at least a multinode of size 3.
  • Case 2: At the beginning round $r$, there is at least one multinode of size 2 and there is no multinode of size more than 2.
Let $v$ be multinode in the ring $\mathcal{R}$ at the beginning round $r$, and $v_{1}$ and $v_{2}$ be two neighbours of node $v$ in the ring $\mathcal{R}$. Since $n\geq 6$, except node $v$, let $w_{1}$ and $w_{2}$ be neighbours of $v_{1}$ and $v_{2}$ respectively on the footprint ring. Agents at node $w_{1}$ can not see the information of node $v_{2}$, $w_{2}$ and $v$ using 1-hop visibility. Similarly, agents at node $w_{2}$ can not see the information of node $v_{1}$, $w_{1}$ and $v$. Therefore, according to $\mathcal{A}$, the decision of the move cycle in each round of agents in node $v_{1}$ is independent of the decision of the move cycle of agents at node $v_{2}.$ At the start of round $r$, if the adversary removes an edge according to the following strategy then at least a multinode remains at the start of the next round.
  • Strategy for case 1: Let $v$ be a multinode of size 3. At the starting of round $r$, the adversary removes edge $(v,\,v_{1})$. In this case, no matter what steps agents follow, there will be a multinode of the size of at least two in the next round.
  • Strategy for case 2: Let $v$ be a multinode of size two. We divide this into three cases, depending on the situation at the beginning of round $r$.
    • Both $v_{1}$ and $v_{2}$ are holes
      If in round $r$, no agent moves to the node $v_{1}$ and $v_{2}$ from their neighbours except node $v$, then there is always a hole present at the start of round $r+1$.
      Else, if in round $r$, agent(s) move to the node $v_{1}$ and $v_{2}$ from their neighbours $w_{1}$ and $w_{2}$ respectively, then there is always a multinode present among $v_{1},v,v_{2}$ at the start of the next round.
      Else, in round $r$, agent(s) move to $v_{1}$ (or $v_{2}$) from its neighbour $w_{1}$ (or $w_{2}$). W.l.o.g., agent(s) decides to move from node $w_{1}$ to node $v_{1}$ in round $r$. In this case, the adversary removes the edge $(v,\,v_{2})$ at the start of round $r$. At the round $r$, agents from node $v$ can not move to the node $v_{2}$ due to the missing edge $(v,\,v_{2})$. They either can stay at node $v$ or move to the node $v_{1}$. In both cases, there is a multinode at the start of the next round. Specifically, if agent(s) at node $v$ decides to move, it creates a multinode at node $v_{1}$; otherwise, node $v$ remains as a multinode at the start of round $r+1$.
    • Either $v_{1}$ or $v_{2}$ is a hole W.l.o.g. let $v_{1}$ be a hole.
      If in round $r$, one agent moves from both neighbours of node $v_{1}$, the adversary does not remove any edge at the start of round $r$. In this case, it creates a multinode at node $v_{1}$ at the start of round $r+1$.
      Else, if in round $r$, agent(s) from one of the nodes $v$, $w_{1}$, say from $w$, decides to move to the node $v_{1}$, the adversary removes edge $(v_{1},\,w)$ at the start of round $r$. In this case, node $v_{1}$ remains a hole at the start of the next round.
      Otherwise, in round $r$, no agent moves to the node $v_{1}$ in the round $r$, then node $v_{1}$ remains hole at the start of round $r+1$.
    • Neither $v_{1}$ nor $v_{2}$ is a hole
      If in round $r$, agents at node $v_{1}$ and $v_{2}$ do not move in round $r$, the adversary does not remove any edge. In this case, one among $v_{1},v,v_{2}$ is a multinode at the start of the next round.
      Otherwise, if in round $r$, agent(s) from node $v_{1}$ and $v_{2}$ move to node $v$, then the adversary does not remove any edge. In this case, there will be a multinode at the node $v$.
      Else, if in round $r$, agent(s) of either $v_{1}$ or $v_{2}$, w.l.o.g. say of $v_{1}$, does not move in round $r$, then the adversary removes the edge $(v,\,v_{2})$. In this case, there will be a multinode at $v$ or $v_{1}$ at the start of the next round.
      Otherwise, if in round $r$, agent(s) from node $v_{1}$ and $v_{2}$ move to node $w_{1}$ and $w_{2}$, respectively, the adversary does not remove any edge. In this case, no matter how agents move from node $v$, it creates a hole or a multinode at the start of round $r+1$.
      Else, in round $r$, agent(s) of either $v_{1}$ or $v_{2}$ move to $v$, say agent(s) of $v_{1}$ move to node $v$. In this case, if the agent from node $v_{2}$ also moves to node $v$, then the adversary does not remove any edge. At the beginning of the next round $r+1$, there is a multinode at node $v$. If the agent at $v_{2}$ decides to stay at node $v_{2}$ in round $r$, then also the adversary does not remove any edge. In this case, there is a multinode at node $v$ or $v_{2}$ at the beginning of the next round. If agent from $v_{2}$ moves toward to $w_{2}$, then the adversary deletes edge $(v_{2},\,w_{2})$. In this case, no matter how the agents move from node $v_{2}$, there is a multinode at node $v_{2}$ or $v$ at the beginning of round $r+1$.
This completes the proof. ∎
Theorem 2.2
(Impossibility without 1-hop visibility) It is impossible to solve Dispersion on TVG with $k=n$ agents considering global communication but without $1$-hop visibility even if the agents are initially co-located and have infinite memory.
Proof
If the agents have zero visibility, then the choice of movement of each agent depends only on the information present at the current node and the information attained via global communication. Since the algorithm is deterministic, the adversary can remove an edge from $G$ based on the movement of the agents. We prove the above theorem by showing the impossibility of achieving dispersion on rings. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movement of agents does not lead to dispersion, the adversary does not need to do anything. If the movement of agents leads to dispersion, the adversary has to intervene. We have the following possible scenarios:
  • If there is a hole, say $v$, in the ring and the agents from both the neighbouring nodes of nodes $v$ are going to enter into the node: In this case, the adversary does not need to do anything as a new multinode is created in the ring.
  • If there is a hole in the ring and the agent from exactly one neighbouring node is going to enter into the hole: In this case, the adversary removes the edge connecting the hole with this neighbouring node (recall that in our model agents at a node $v$ can not sense whether an edge associated with node $v$ is missing or not). Therefore, the movement of the agent becomes unsuccessful in this round. Thus, the hole remains, and dispersion is not achieved.
The adversary can prevent the agents from achieving dispersion in any given round $r$ using the above-mentioned strategies. ∎
Remark 1. From [7], the time lower bound is $\Omega(k)$ to solve dispersion on static graphs for $k\leq n$. From [2], the memory lower bound is $\Omega(\log k)$ per agent to solve dispersion on static graphs for $k\leq n$. We are working on a dynamic graph model which is weaker than a static graph model. Therefore, the lower bound for time and memory per agent to solve Dispersion on TVG for $k\leq n$ is $\Omega(k)$ and $\Omega(\log k)$, respectively. Due to Theorem 2.1 and 2.2 to solve Dispersion on TVG $\mathcal{G}$ for $k\leq n$, agents need 1-hop visibility and global communication. In [10], when agents are equipped with 1-hop visibility and global communication, authors provide an algorithm to solve dispersion in $O(k)$ time and $O(\log k)$ memory per agent in their dynamic graph model. Therefore, we can use the same algorithm and solve Dispersion on TVG for $k\leq n$ agents in $\Theta(k)$ time and $\Theta(\log k)$ memory per agent as TVG exhibits a stronger graph model.
3 The Dispersion on General Graphs
In this section, we introduce an algorithm that solves the Dispersion on TVG when there are $n+1$ agents co-located at some node in a general graph. We use the idea of depth-first search (DFS) traversal by mobile agents that also solves the dispersion on static graphs [2]. For the sake of completeness, let us recall the idea of DFS traversal. At each node $v$, an agent chooses an unexplored port $p$ and traverses through it: (i) if it reaches an unvisited node $u$ (agents can understand this by the fact that there is no agent settled at $u$), then one agent settles at $u$ and stores port $p^{\prime}$ (through which it entered node $u$) as its parent port. Now, the unsettled agent tries to move through port $(p^{\prime}+1)$ mod $\delta_{u}$ in the explore state, (ii) else, if the agent reaches an already visited node (it can understand this because there is a settled agent present at that node), the agent returns to node $v$ in the backtrack state. If an unsettled agent reaches node $v$ in the backtrack state, then it computes $q=(p+1)$ mod $\delta_{v}$. If port $q$ is unexplored, it moves through port $q$ in the explore state. Else, if port $q$ equals the parent port of $v$, then it moves through port $q$ in the backtrack state. This algorithm takes $4m$ rounds to run on any graph, where $m$ is the number of edges in that graph. The information required for an agent at each node to run this algorithm can be stored by the settled agent that is already present at that node.
Recall that, according to our model, an agent does not know if any associated edge w.r.t. its current position is missing or not. This makes the adversary powerful as it can precompute the agents’ decision at a node and accordingly can delete an associated edge at the start of that round. If the group of unsettled agents decides to move through port $p$ according to our algorithm in round $r$, the adversary can delete the edge corresponding to port $p$ in round $r$. Thus preventing the group of agents from moving in round $r$.
High-level idea: No group of agents moves indefinitely through which its movement was unsuccessful. Initially, there is only one group as the configuration is rooted. Whenever the group has an unsuccessful move, the agents divide into two groups, namely $G_{1}$ and $G_{2}$. These two groups run their dispersion algorithm via DFS separately. A missing edge can make the movement unsuccessful for both groups. On unsuccessful movement, $G_{1}$ never deviates from its DFS traversal path, whereas $G_{2}$ deviates after some finite waiting period. This keeps the algorithm running and either $G_{2}$ gets dispersed or $G_{1}$ can move one hop according to its DFS traversal. If $G_{2}$ gets dispersed, $G_{1}$ can understand it within some finite rounds and gets divided into two groups further. Else, if $G_{1}$ gets dispersed by moving one hop again and again, $G_{2}$ can understand it within some finite rounds and gets divided into two groups further. We ensure that none of $G_{1}$ (or $G_{2}$) divides further unless $G_{2}$ (or $G_{1}$) is already dispersed. This keeps the total number of groups bounded by 2. As the settled agents keep the information required for the DFS traversal of all the groups, it helps the agents store the information using $O(\log n)$ memory. As we have different algorithms for the groups and different criteria for the groups to be divided, while dividing a group into two, each new group is given a unique label so that $G_{1}\,(G_{2})$ can execute their respective algorithms.
3.1 Algorithm for general graph
In this section, we provide a detailed description of the algorithm for dispersion on a time-varying graph $\mathcal{G}$ with at most one dynamic edge. In our algorithm, unsettled agents may get divided into two groups, we call them $Group_{1}$, and $Group_{2}$ and denote them by $G_{1}$, and $G_{2}$ respectively. At any particular round, there can be no more than two groups. However, over time, if all the unsettled agents of one group get settled, the other group can be divided into two groups again. So, if we look throughout the run of the algorithm, there may exist several groups. To identify these groups, we provide unique labelling to each group whenever a group gets generated and store it in a variable. Now we provide the parameters maintained by each agent $a_{i}$.
  • $\bm{a_{i}.settled}$: It is a binary variable and indicates whether $a_{i}$ is settled at the current node or not. If $a_{i}.settled=0$ then it is not settled whereas $a_{i}.settled=1$ denotes that agent $a_{i}$ is settled. Initially $a_{i}.settled=0$.
  • $\bm{a_{i}.prt\_in}$: It stores the port through which an agent $a_{i}$ enters into the current node. Initially, $a_{i}.prt\_in=-1$.
  • $\bm{a_{i}.prt\_out}$: It stores the port through which an agent $a_{i}$ exits from the current node. Initially, $a_{i}.prt\_out=-1$.
  • $\bm{a_{i}.state}$: It denotes the state of an agent $a_{i}$ that can be either explore or backtrack. Initially $a_{i}.state=explore$.
  • $\bm{a_{i}.dfs\_label}$: It stores the number of new DFS traversals started by agent $a_{i}$. Initially $a_{i}.dfs\_label=1$.
  • $\bm{a_{i}.success}$: When an agent $a_{i}$ attempts to move from a node $u$ to $v$ via the edge $(u,\,v)$, in the round $t$, it is either successful or unsuccessful. If the agent reaches the node $v$ in the round $t+1$, it is a successful attempt, thus, $a_{i}.success=True$; otherwise, $a_{i}.success=False$. Each agent $a_{i}$ initially has $a_{i}.success=True$.
  • $\bm{a_{i}.skip}$: If agent $a_{i}$ skips a port, then it stores in this parameter. Initially, $a_{i}.skip=-1$
  • $\bm{a_{i}.divide}$: Initially, all agents are together and each agent has $a_{i}.divide=0$. If and when this group of agents divides into two groups for the first time, each agent sets $a_{i}.divide=1$. There is no change in the values of this variable thereafter.
  • $\bm{a_{i}.grp\_label}$: It stores the ID of the group to which agent $a_{i}$ belongs. Initially $a_{i}.grp\_label=10$. If $a_{i}.grp\_label$ ends with 0, it implies $a_{i}$ is a part of $G_{1}$. Else, if $a_{i}.grp\_label$ ends with 1, it implies $a_{i}$ is a part of $G_{2}$. Later we elaborate on how this variable gets updated.
  • $\bm{a_{i}.count_{1}}$: If agent $a_{i}\in G_{1}$, then it stores the number of consecutive odd rounds with $a_{i}.success=False$.
  • $\bm{a_{i}.count_{2}}$: If agent $a_{i}\in G_{2}$ then it stores the number of consecutive odd rounds with $a_{i}.success=False$.
  • $\bm{a_{i}.count_{3}}$: If agent $a_{i}\in G_{2}$, then it stores the number of new DFS traversals it started from the time $a_{i}$ became a part of this group.
Let $a_{i}$ be a settled agent at node $v$. Any unsettled agent belongs to either $G_{1}$ or $G_{2}$. When an unsettled agent $a_{j}$ while doing its movement (along with other unsettled agents, if any, in its group) according to our algorithm, reaches a node $v$ where an agent $a_{i}$ is already settled, $a_{j}$ needs some information regarding $v$ from $a_{i}$ to continue its traversal. In this regard, each settled agent maintains two parameters, one w.r.t. each of the groups as mentioned below.
[…]
Suppose $n+1$ agents are co-located at a node say $v$ of the graph (rooted configuration). Initially, for each agent $a_{i}$, $a_{i}.grp\_label=10$. Therefore, all agents are a part of $G_{1}$. According to our model, an agent cannot sense the presence of a missing edge unless it attempts to move through it. To tackle this, we take two consecutive rounds to perform one edge traversal. In every odd round, agents move, and in the next even round, it realizes the movement is successful or unsuccessful. Precisely, say an agent $a_{i}$ attempts to move in an odd round $r$. If its movement is successful, it updates its parameter $a_{i}.success$ to $True$ in the even round $r+1$. Otherwise, it updates $a_{i}.success=False$ in the even round $r+1$. Now we proceed with a detailed description of our algorithm.
Algorithm for unsettled agent $\bm{a_{i}}$ of $\bm{G_{1}}$ at odd round: In any odd round, let agent $a_{i}\in G_{1}$ be at node $u$ and $a_{i}.settled=0$. It follows the following steps.
  • $\bm{a_{i}.divide=0}$: Suppose, agent $a_{i}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. Else if agent $a_{i}$ is not the minimum ID, then it moves according to the DFS traversal algorithm. If agent $a_{i}$ reaches a node $v$ and finds an agent with $a_{j}$ with $a_{j}.settled=1$, then it $a_{i}$ does not settle at node $v$ and moves according to the DFS traversal algorithm. In this way, the agent $a_{i}\in G_{1}$ continues the DFS traversal of the graph unless $a_{i}$ encounters a missing edge for the first time. In particular, at some odd round $r$, agent $a_{i}$ moves via an edge $e$, and the movement is not successful at the end of round $r$. This is understood by agent $a_{i}$ in the round $r+1$. In round $r+2$, agent $a_{i}$ updates $divide$ by 1 and divides into two groups using the $Group\_divide$ procedure as follows.
[…]
  • $\bm{a_{i}.divide=1}$: Suppose, agent $a_{i}\in G_{1}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{2}$ at node $w$, then $a_{i}$ updates $a_{i}^{u}(G_{2})$ to $(a_{j}.prt\_in,a_{j}.grp\_label,a_{j}.dfs\_label)$. If $a_{i}\in G_{1}$ reaches a node $w$ and find agent $a_{j}$ with $a_{j}.settled=1$, then it does not settle at the node $v$ and it can understand whether node $w$ is the new node according to the current DFS or previously visited node by comparing $label$, $dfs\_label$ components of $a_{j}^{w}(G_{1})$ with $a_{i}.grp\_label$, $a_{i}.dfs\_label$ respectively. If both matches then node $w$ is visited previously in the current DFS traversal of agent $a_{i}$. Otherwise, if any of those two components do not match, then agent $a_{i}\in G_{1}$ treats $w$ as an unvisited node w.r.t. the current DFS. Round $r+2$ onwards, $a_{i}$ updates its $count_{1}$ parameter as follows. At odd round $r_{1}\geq r+2$, if it finds $a_{i}.success=True$, i.e. it realized a successful movement in the even round $r_{1}-1$, agent $a_{i}$ updates $a_{i}.count_{1}=0$; else if $a_{i}.success=False$, agent $a_{i}$ updates $a_{i}.count_{1}=a_{i}.count_{1}+1$. If at some odd round, $a_{i}$ finds $a_{i}.count_{1}=16n^{2}$, then unsettled agents at node $u$ divide into two groups again using aforementioned $Group\_divide$ procedure. That is, $G_{1}$ divides into two groups only after getting stuck at a node for a particular port for $16n^{2}$ many odd rounds.
Algorithm for unsettled agent $\bm{a_{i}}$ of $\bm{G_{2}}$ at odd round: In any odd round, let agent $a_{i}\in G_{2}$ be at node $u$ and $a_{i}.settled=0$. It follows the following steps.
  • (1)
    $\bm{a_{i}.state=explore}$:
    • (A)
      $\bm{a_{i}.success=True}$: It updates $a_{i}.count_{2}=0$ and does the following.
      • (a)
        Node $\bm{u}$ has no agent with $\bm{setlled=1}$: If $a_{i}$ is the minimum ID at node $u$, it settles at node $u$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ and $a_{i}^{u}(G_{2})$ to $(a_{i}.prt\_in,a_{i}.grp\_label,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{1}$ at node $u$, then $a_{i}$ updates $a_{i}^{u}(G_{1})$ to $(a_{j}.prt\_in,\\ a_{j}.grp\_label,a_{j}.dfs\_label)$. If $a_{i}$ is not the minimum ID, then it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
      • (b)
        Node $\bm{u}$ has an agent with $\bm{setlled=1}$: Let $a_{j}$ be at node $u$ with $a_{j}.settled=1$. If $a_{i}.grp\_label$, $a_{i}.dfs\_label$ matches $label$, $dfs\_label$ components of $a_{j}^{u}(G_{2})$ respectively, then it is a previously visited node. In this case $a_{i}$ set $a_{i}.state=backtrack$, $a_{i}.prt\_out=a_{i}.prt\_in$, and it moves through $a_{i}.prt\_out$. Otherwise, it is an unvisited node w.r.t. the current DFS of $a_{i}$. In this case, it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
    • (B)
      $\bm{a_{i}.success=False}$: It does the following.
      • (a)
        There is an agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: Let $a_{j}$ be settled at node $u$, $b_{i}\in G_{1}$ be at node $u$, and $b_{i}.settled=0$. In this case, there are two cases.
        • (i)
          If $a_{i}.prt\_out=b_{i}.prt\_out$ (observe that based on communication, agent $a_{i}$ can compute what agent $b_{i}$ computes, and vice versa), then there are two cases possible.
          ($\alpha$) $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is $\bm{-1}$: If $a_{i}.prt\_out=\delta_{u}-1$, then agent $a_{i}$ sets $a_{i}.state=explore$, stores $a_{i}.skip=-1$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=0$, and moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out\neq\delta_{u}-1$, then it sets $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{u}$ and sets $a_{i}.count_{2}=0$ and move through $a_{i}.prt\_out$.
          ($\beta$) $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is not $\bm{-1}$: It sets $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{u}$ and sets $a_{i}.count_{2}=0$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
        • (ii)
          If $a_{i}.prt\_out\neq b_{i}.prt\_out$, it updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. In this case, agent $a_{i}$ sets $a_{i}.state=explore$, stores $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Note that $a_{i}.count_{3}$ always remains less than $4n^{2}$ in this case as justified in the correctness.
      • (b)
        There is no agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: It updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Else, if $a_{i}.count_{3}=4n^{2}$, it divides into two groups using $Group\_divide$ procedure mentioned in the algorithm for unsettled agents of $G_{1}$. Note that if $a_{i}.count_{3}=4n^{2}$, then $G_{1}$ is already dispersed as justified in the correctness.
  • (2)
    $\bm{a_{i}.state=backtrack}$:
    • (A)
      If $\bm{a_{i}.success=True}:$ It updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$ and $a_{i}.count_{2}=0$. Let $a_{j}$ be the agent at node $u$ with $a_{j}.settled=1$. Based on $parent$ component of $a_{j}^{u}(G_{2})$, there are two cases.
      • (a)
        $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is $\bm{-1}$: If agent $a_{i}$ finds $parent$ component of $a_{j}^{u}(G_{2})$ to be $-1$, then node $u$ is the root node of the current DFS traversal of agent $a_{i}$. In this case, if $a_{i}.prt\_out=$ minimum available port except for $a_{i}.skip$, then it sets $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, and moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out\neq$ minimum available port except for $a_{i}.skip$, then it does the following.
        If $a_{i}.prt\_out=a_{i}.skip$, then $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{v}$. If $a_{i}.prt\_out=$ minimum available port except for $a_{i}.skip$, then it sets $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, and moves through $a_{i}.prt\_out$. Else if, it sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
        If $a_{i}.prt\_out\neq a_{i}.skip$, it sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
      • (b)
        $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is not $\bm{-1}$: If agent $a_{i}$ does not find $parent$ component of $a_{j}^{u}(G_{2})$ to be $-1$, then node $u$ is not the root node of the current DFS traversal of agent $a_{i}$. If $a_{i}.prt\_out$ matches with $parent$ component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out$ does not match with $parent$ component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
    • (B)
      If $\bm{a_{i}.success=False}:$ Two cases are possible.
      • (a)
        There is an agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: Let $a_{j}$ be settled at node $u$, $b_{i}\in G_{1}$ be at node $u$, and $b_{i}.settled=0$. In this case, there are two cases.
        • (i)
          If $a_{i}.prt\_out=b_{i}.prt\_out$ (observe that based on communication, agent $a_{i}$ can compute what agent $b_{i}$ computes, and vice versa), then agent $a_{i}$ does the following. It updates $a_{i}.count_{2}=0$ and sets $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$,
          $a_{i}.dfs\_label=a_{i}.dfs\_label+1$. Agent $a_{i}$ moves through $a_{i}.prt\_out$.
        • (ii)
          If $a_{i}.prt\_out\neq b_{i}.prt\_out$, it updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.port=a_{i}.prt\_out$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=0$, $a_{i}.skip=-1$, and moves through $a_{i}.prt\_out$. Note that $a_{i}.count_{3}$ always remains less than $4n^{2}$ in this case as justified in the correctness.
      • (b)
        There is no agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: It updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Else, if $a_{i}.count_{3}=4n^{2}$, it divides into two groups using $Group\_divide$ procedure mentioned in the algorithm for unsettled agents of $G_{1}$.
Algorithm for unsettled agent $\bm{a_{i}}$ of at even round: At even round, each unsettled agent $a_{i}$ of $G_{1}$ or $G_{2}$ updates $a_{i}.success$ parameter. If their movement in the previous rounds is successful, agent $a_{i}$ updates their $a_{i}.success=True$. Otherwise, it updates their parameter $a_{i}.success=False$.
Algorithm for settled agents: Let $a_{i}$ be at node $u$ with $a_{i}.settled=1$. It has two parameter $a_{i}^{u}(G_{1})$ and $a_{i}^{u}(G_{2})$. At each odd round based on the communication, it computes the decision of $G_{1}$ and $G_{2}$, and it stores $G_{1}$ information in $a_{i}^{u}(G_{1})$ and $G_{2}$ information in $a_{i}^{u}(G_{2})$. In move, it does not do anything. At even round, it does not do anything.
The pseudocode of our algorithm, Dispersion in 1-TVG with $n+1$ agents, is provided in Algorithm 1.
1 while True do
2       $a_{i}.r=a_{i}.r+1$
3       if $a_{i}$ belongs to $G_{1}$ and $a_{i}.settled=0$ then
4             call Algorithm 3
5      else if $a_{i}$ belongs to $G_{2}$ and $a_{i}.settled=0$ then
6             call Algorithm 2
7      else if $a_{i}.settled=1$, there is $a_{j}$ from $G_{1}$ at the same node and $a_{j}.settled=0$ then
8             call Algorithm 7
9      else if $a_{i}.settled=1$, there is $a_{j}$ from $G_{2}$ at the same node and $a_{j}.settled=0$ then
10             call Algorithm 8
Algorithm 1 Dispersion in $1$-TVG with $n+1$ agents
1 if $a_{.}state=explore$ then
2       call Algorithm 4
3else
4       call Algorithm 5
Algorithm 2 Movement of agent $a_{i}\in G_{2}$ at $v$ with $a_{i}.settled=0$
1 if $a_{i}.state=explore$ then
2       if $r$ mod 2 =1 then
3             if $a_{i}.success=True$ then
4                   set $a_{i}.count_{1}=0$
5                   if there is no agent at node $v$ with $settled=1$ then
6                         if $a_{i}.ID$ is the minimum from all the unsettled agents present at the current node then
7                               set $a_{i}.settled=1$
8                               update $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$
9                               if there is an agent $a_{j}$ from $G_{2}$ with $a_{j}.settled=0$ then
10                                     update $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
11                        else
12                               set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
13                               if $a_{i}.prt\_out=a_{i}.prt\_in$ then
14                                     set $a_{i}.state=backtrack$
15                                     move through $a_{i}.prt\_out$
16                              else
17                                     move through $a_{i}.prt\_out$
18                  else if there is an agent $r_{j}$ at node $v$ with $settled=1$ then
19                         $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{v}$
20                         if $label=a_{i}.grp\_label$, where $label$ is the component of $r_{j}^{v}(G_{1})$ then
21                               $a_{i}.state=backtrack$
22                               move through $a_{i}.prt\_in$
23                        else if $label\neq a_{i}.grp\_label$, where $label$ is the component of $r_{j}^{v}(G_{1})$ then
24                               move through $a_{i}.prt\_out$
25            else if $a_{i}.success=False$ then
26                   if $a_{i}.divide=0$ then
27                         set $a_{i}.divide=1$
28                         call $Group\_divide()$ (Algorithm 6)
29
30                  else if $a_{i}.divide=1$ then
31                         update $a_{i}.count_{1}=a_{i}.count_{1}+1$
32                         if $a_{i}.count_{1}=16n^{2}$ then
33                               call $Group\_divide()$ (Algorithm 6)
34
35                        else if $a_{i}.count_{1}<16n^{2}$ then
36                               move through $a_{i}.prt\_out$
37      else if $r$ mod 2=0 then
38             if the movement of agent $a_{i}$ in round $r-1$ is successful then
39                   set $a_{i}.success=True$
40            else
41                   set $a_{i}.success=False$
42else if $a_{i}.state=backtrack$ then
43       if $r$ mod 2=1 then
44             if $a_{i}.success=True$ then
45                   set $a_{i}.count_{1}=0$
46                   if $parent$ component of $r_{j}^{v}(G_{1})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
47                         set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
48                         set $a_{i}.state=explore$
49                         move through $a_{i}.prt\_out$
50                  else if $parent$ component of $r_{j}^{v}(G_{1})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
51
52                        set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
53                         if $a_{i}.prt\_out=parent$ and $parent$ is the component of $r_{j}^{v}(G_{1})$, where $r_{j}$ is the settled agent at node $v$ then
54                               move through $a_{i}.prt\_out$
55                        else
56                               set $a_{i}.state=explore$
57                               move through $a_{i}.prt\_out$
58            else if $a_{i}.success=False$ then
59                   if $a_{i}.divide=0$ then
60                         set $a_{i}.divide=1$
61                         call $Group\_divide()$ (Algorithm 6)
62
63                  else if $a_{i}.divide=1$ then
64                         update $a_{i}.count_{1}=a_{i}.count_{1}+1$
65                         if $a_{i}.count_{1}=16n^{2}$ then
66                               call $Group\_divide()$ (Algorithm 6)
67
68                        else if $a_{i}.count_{1}<16n^{2}$ then
69                               move through $a_{i}.prt\_out$
70      if $r$ mod 2=0 then
71             if the movement of agent $a_{i}$ in round $r-1$ is successful then
72                   set $a_{i}.success=True$
73            else
74                   set $a_{i}.success=False$
Algorithm 3 Movement of agent $a_{i}\in G_{1}$ at $v$ with $a_{i}.settled=0$
1 if $r$ mod 2 =1 then
2       if $a_{i}.success=True$ then
3             set $a_{i}.count_{2}=0$
4             if there is no agent at node $v$ with $settled=1$ then
5                   if $a_{i}.ID$ is the minimum from all the unsettled agents present at node $v$ then
6                         set $a_{i}.settled=1$
7                         update $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$
8                         if there is an agent $a_{j}$ from $G_{1}$ with $a_{j}.settled=0$ at node $v$ then
9                               update $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
10                  else
11                         set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
12                         if $a_{i}.prt\_out=a_{i}.prt\_in$ then
13                               set $a_{i}.state=backtrack$
14                        move through $a_{i}.prt\_out$
15            else if there is an agent $r_{j}$ at node $v$ with $r_{j}.settled=1$ then
16                   if $a_{i}.grp\_label=label$ and $a_{i}.dfs\_label=dfs\_label$, where $label$ and $dfs\_label$ are the component from $r_{j}^{v}(G_{2})$ then
17                         set $a_{i}.state=backtrack$
18                         set $a_{i}.prt\_out=a_{i}.prt\_in$
19                         move through $a_{i}.prt\_out$
20                  else if $a_{i}.grp\_label\neq label$ or $a_{i}.dfs\_label\neq dfs\_label$, where $label$ and $dfs\_label$ are the component from $r_{j}^{v}(G_{2})$ then
21                         set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
22                         if $a_{i}.prt\_out=a_{i}.prt\_in$ then
23                               set $a_{i}.state=backtrack$
24                               move through $a_{i}.prt\_out$
25                        else
26                               move through $a_{i}.prt\_out$
27      else if $a_{i}.success=False$ then
28             if  $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out=a_{i}.prt\_out$ then
29                   if $parent$ component of $r_{j}^{v}(G_{2})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
30                         if $a_{i}.prt\_out=\delta_{v}-1$ then
31                               $a_{i}.count_{2}=0$
32                               $a_{i}.skip=-1$
33                               $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
34                               $a_{i}.prt\_out=0$
35                               move through $a_{i}.prt\_out$
36                        else if $a_{i}.prt\_out\neq\delta_{v}-1$ then
37                               $a_{i}.count_{2}=0$
38                               $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{v}$
39                               move through $a_{i}.prt\_out$
40                  else if $parent$ component of $r_{j}^{v}(G_{2})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
41                         $a_{i}.count_{2}=0$
42                         $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{v}$
43                         if $a_{i}.prt\_out=a_{i}.prt\_in$ then
44                               $a_{i}.state=backtrack$
45                               move through $a_{i}.prt\_out$
46                        else if $a_{i}.prt\_out\neq a_{i}.prt\_in$ then
47                               move through $a_{i}.prt\_out$
48            else if $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out\neq a_{i}.prt\_out$ or there is no agent from $G_{1}$ present at node $v$ then
49                   set $a_{i}.count_{2}=a_{i}.count_{2}+1$
50                   if $a_{i}.count_{2}<4n^{2}$ then
51                         move through $a_{i}.prt\_out$
52                  else if $a_{i}.count_{2}=4n^{2}$ then
53                         $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$
54                         if $a_{i}.count_{3}<4n^{2}$ then
55                               set $a_{i}.state=explore$
56                               set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
57                               set $a_{i}.skip=a_{i}.prt\_out$
58                               set $a_{i}.prt\_out$ as the minimum port available at the current node except $a_{i}.skip$
59                               move through $a_{i}.prt\_out$
60                        else
61                               call $Group\_divide()$ (Algorithm 6)
62
63else if $r$ mod 2=0 then
64       if the movement of agent $a_{i}$ in round $r-1$ is successful then
65             set $a_{i}.success=True$
66      else
67             set $a_{i}.success=False$
Algorithm 4 $a_{i}.state=explore$
1 if $r$ mod 2 =1 then
2       if $a_{i}.success=True$ then
3             set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
4             set $a_{i}.count_{2}=0$
5             if  $parent$ component of $r_{j}^{v}(G_{2})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
6                   if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
7                         $a_{i}.state=explore$
8                         $a_{i}.skip=-1$
9                         $a_{i}.prt\_out=0$
10                         $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
11                         move through $a_{i}.prt\_out$
12                  else if $a_{i}.prt\_out\neq$ minimum available port except $a_{i}.skip$ then
13                         if $a_{i}.prt\_out=a_{i}.skip$ then
14                               $a_{i}.prt\_out=(a_{i}.prt\_out+1)\mod\delta_{v}$
15                               if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
16                                     $a_{i}.state=explore$
17                                     $a_{i}.skip=-1$
18                                     $a_{i}.prt\_out=0$
19                                     $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
20                                     move through $a_{i}.prt\_out$
21                              else if $a_{i}.prt\_out\neq$ minimum available port except $a_{i}.skip$ then
22                                     $a_{i}.state=explore$
23                                     move through $a_{i}.prt\_out$
24                        else if $a_{i}.prt\_out\neq a_{i}.skip$ then
25                               $a_{i}.state=explore$
26                               move through $a_{i}.prt\_out$
27            else if $parent$ component of $r_{j}^{v}(G_{2})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
28                   if $a_{i}.prt\_out=parent$, where $parent$ is the component of $r_{j}^{v}(G_{2})$ then
29                         move through $a_{i}.prt\_out$
30                  else if $a_{i}.prt\_out\neq parent$, where $parent$ is the component of $r_{j}^{v}(G_{2})$ then
31                         $a_{i}.state=explore$
32                         move through $a_{i}.prt\_out$
33      else if $a_{i}.success=False$ then
34
35            if $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out=a_{i}.prt\_out$ then
36                   $a_{i}.count_{2}=0$
37                   set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
38                   set $a_{i}.state=explore$
39                   $a_{i}.skip=a_{i}.prt\_out$
40                   set $a_{i}.prt\_out$ as the minimum port available at the current node except $a_{i}.skip$
41                   move through $a_{i}.prt\_out$
42            else if $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out\neq a_{i}.prt\_out$ or there is no agent from $G_{1}$ present at node $v$ then
43                   $a_{i}.count_{2}=a_{i}.count_{2}+1$
44                   if $a_{i}.count_{2}=4n^{2}$ then
45                         set $a_{i}.count_{2}=0$, and $a_{i}.count_{3}=a_{i}.count_{3}+1$
46                         if $a_{i}.count_{3}<4n^{2}$ then
47                               set $a_{i}.state=explore$
48                               set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
49                               set $a_{i}.skip=a_{i}.prt\_out$
50                               set $a_{i}.prt\_out$ as the minimum port available at the current node except $a_{i}.prt\_out$
51                               move through $a_{i}.prt\_out$
52                        else
53                               call $Group\_divide()$ (Algorithm 6)
54
55                  else if $a_{i}.count_{2}<4n^{2}$ then
56                         $a_{i}.count_{2}=a_{i}.count_{2}+1$
57                         move through $a_{i}.prt\_out$
58else if $r$ mod 2 = 0 then
59       if the movement of agent $a_{i}$ in round $r-1$ is successful then
60             set $a_{i}.success=True$
61      else
62             set $a_{i}.success=False$
Algorithm 5 $a_{i}.state=backtrack$
1 let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs
2 if $i\leq\lceil{\frac{x}{2}}\rceil$ then
[…]
4       $a_{i}.prt\_out=0$
5       set $a_{i}.state=explore$
6       $a_{i}.skip=-1$
7       set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
8       move through $a_{i}.prt\_out$
9else
[…]
11       $a_{i}.prt\_out=0$ set $a_{i}.state=explore$
12       $a_{i}.skip=-1$
13       set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
14       move through $a_{i}.prt\_out$
Algorithm 6 $Group\_divide()$
1 if $r$ mod 2=1 then
2       if unsettled agent $a_{j}$ from $G_{1}$ and $a_{j}.success=True$ then
3             if unsettled agents $a_{j}$ from $G_{1}$ is present at node $v$ with $a_{j}.state=backtrack$ then
4                   It does not do anything.
5            if unsettled agents $a_{j}$ from $G_{1}$ is present at node $v$ with $a_{j}.state=explore$ then
6                   if $label$ component of $a_{i}^{v}(G_{1})$ does not match with $a_{j}.grp\_label$ then
7                         set $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
8                  else if $label$ component of $a_{i}^{v}(G_{1})$ matches with $a_{j}.grp\_label$ then
9                         It does not do anything
10      else if unsettled agent $a_{j}$ from $G_{1}$ and $a_{j}.success=False$ then
11             if unsettled agents $a_{j}$ from $G_{1}$ is present at node $v$ with $a_{j}.state=backtrack$ or $explore$ then
12                   if $a_{j}.count_{1}+1<16n^{2}$ then
13                         It does not do anything.
14                  else if $a_{j}.count_{1}+1=16n^{2}$ then
15                         let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs. Let $L=a_{i}.grp\_label=y$, $i\in[1,\,x]$
16                         if $i\leq\lceil{\frac{x}{2}}\rceil$ then
17                               set $a_{i}^{v}(G_{1}).(parent,label,dfs\_label)=(-1,\,L.0,\,a_{j}.dfs\_label+1)$
18                        else
19                               set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,L.1,\,a_{j}.dfs\_label+1)$
20
21else if $r$ mod 2=0 then
22       It does not do anything.
Algorithm 7 Algorithm for settled agent $a_{i}$ at node $v$ and at least one unsettled agent $a_{j}$ from $G_{1}$ present at node $v$
1 if $r$ mod 2=1 then
2       if unsettled agent $a_{j}$ from $G_{2}$ and $a_{j}.success=True$ then
3             if $a_{j}.state=backtrack$ then
4                   if $parent$ component of $a_{i}^{v}(G_{2})$ is $-1$ then
5                         if $(a_{j}.prt\_out+1)$ mod $\delta_{v}=$ minimum available port except for $a_{j}.skip$ then
6                               set $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,a_{j}.grp\_label,\,a_{j}.dfs\_label+1)$
7                        else if $(a_{j}.prt\_out+1)$ mod $\delta_{v}\neq$ minimum available port except for $a_{j}.skip$ then
8                               if $(a_{j}.prt\_out+1)$ mod $\delta_{v}=a_{j}.skip$ then
9                                     if $(a_{j}.prt\_out+2)=\delta_{v}-1$ then
10                                           set $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,a_{j}.grp\_label,\,a_{j}.dfs\_label+1)$
11                                    else if $(a_{j}.prt\_out+2)\neq\delta_{v}-1$ then
12                                           It does not do anything
13                              else if $(a_{j}.prt\_out+1)$ mod $\delta_{v}\neq a_{j}.skip$ then
14                                     It does not do anything
15                  else if $parent$ component of $a_{i}^{v}(G_{2})$ is not -1 then
16                         It does not do anything
17            if $a_{j}.state=explore$ then
18                   if $label$ and $dfs\_label$ components from $a_{i}^{v}(G_{2})$ do not match with $a_{j}.grp\_label$ and $a_{j}.dfs\_label$ then
19                         set $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
20                  if $label$ and $dfs\_label$ components from $a_{i}^{v}(G_{2})$ match with $a_{j}.grp\_label$ and $a_{j}.dfs\_label$ then
21                         It does not do anything
22      else if unsettled agent $a_{j}$ from $G_{2}$ and $a_{j}.success=False$ then
23             if unsettled agents $b_{i}$ from $G_{1}$ is present at node $v$ then
24                   if $b_{i}.prt\_out=a_{j}.prt\_out$ and $a_{j}.state=backtrack$ then
25                         $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,a_{i}.grp\_label\,a_{i}.dfs\_label+1)$
26                  else if $b_{i}.prt\_out\neq a_{j}.prt\_out$ and $a_{j}.state=backtrack$ or there is no agent from $G_{1}$ then
27                         if $a_{j}.count_{2}+1=4n^{2}$ then
28                               if $a_{j}.count_{3}+1<4n^{2}$ then
29                                     set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,a_{j}.grp\_label,a_{j}.dfs\_label+1)$
30                              else if $a_{j}.count_{3}+1=4n^{2}$ then
31                                     let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs. Let $L=a_{j}.grp\_label$, $j\in[1,\,x]$
32                                     if $j\leq\lceil{\frac{x}{2}}\rceil$ then
33                                           set $a_{i}^{v}(G_{1}).(parent,label,dfs\_label)=(-1,\,L.0,\,a_{j}.dfs\_label+1)$
34                                    else
35                                           set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,L.1,\,a_{j}.dfs\_label+1)$
36                        else if $a_{j}.count_{2}+1<4n^{2}$ then
37                               It does not do anything
38                  else if $b_{i}.prt\_out=a_{j}.prt\_out$ and $a_{j}.state=explore$ then
39                         if $parent$ component of $a_{i}^{v}(G_{2})$ is $-1$ then
40                               if $a_{i}.prt\_out+1=\delta_{v}-1$ then
41                                     $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,a_{i}.grp\_label,\,a_{j}.dfs\_label+1)$
42                              else if $a_{i}.prt\_out+1\neq\delta_{v}-1$ then
43                                     It does not do anything
44                  else if $b_{i}.prt\_out\neq a_{j}.prt\_out$ and $a_{j}.state=explore$ or there is no agent from $G_{1}$ then
45                         if $a_{j}.count_{2}+1=4n^{2}$ then
46                               if $a_{j}.count_{3}+1<4n^{2}$ then
47                                     set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,a_{j}.grp\_label,a_{j}.dfs\_label+1)$
48                              else
49                                     let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs. Let $L=a_{j}.grp\_label$, $j\in[1,\,x]$
50                                     if $j\leq\lceil{\frac{x}{2}}\rceil$ then
51                                           set $a_{i}^{v}(G_{1}).(parent,label,dfs\_label)=(-1,\,L.0,\,a_{j}.dfs\_label+1)$
52                                    else
53                                           set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,L.1,\,a_{j}.dfs\_label+1)$
54
55                        else if $a_{j}.count_{2}+1<4n^{2}$ then
56                               It does not do anything
57else if $r$ mod 2=0 then
58       It does not do anything.
Algorithm 8 Algorithm for settled agent $a_{i}$ at node $v$ and at least one unsettled agent $a_{j}$ from $G_{2}$ is present at node $v$
3.2 Correctness and analysis of algorithm
Let $m$ be the number of edges in $G$. At the beginning of the algorithm, $n+1$ agents are co-located at a node of $G$. Therefore, all agents can learn the value of $n$. Initially, each agent $a_{i}$ is part of $G_{1}$ as $a_{i}.grp\_label=10$. Assume that the adversary does not remove an edge in the first $4m$ rounds. In this case, all agents reach the dispersion configuration because this is nothing but the dispersion algorithm using the DFS traversal of [2]. Hence, all agents reach dispersion configuration in the first $4m$ rounds.
Let in round $r<4m$, all agents in $G_{1}$ be at node $u$ and want to go through edge $e=(u,\,v)$. At the round $r$, if the adversary removes an edge $e$, then the movement of agents in $G_{1}$ becomes unsuccessful. In this case, we divide the group $G_{1}$ into two new groups $G_{1}$ and $G_{2}$. And, both the groups consider the current node $u$ as the root node for the new DFS traversal and restart the algorithm for DFS via port 0. Let $r^{\prime}\geq r$ be round when agents in $G_{1}$ and $G_{2}$ are at node $u^{\prime}$ and want to go through edge $e^{\prime}=(u^{\prime},\,v^{\prime})$ via port $p$, and their movement becomes unsuccessful due to the deletion of the edge $e^{\prime}$ by the adversary. We have the following claim.
Claim 1: If edge $e^{\prime}$ does not appear within next $16n^{2}$ odd rounds after it is deleted by the adversary in round $r^{\prime}$, then agents in $G_{2}$ get dispersed in those $16n^{2}$ rounds.
Proof
In round $r^{\prime}$, for $a_{i}\in G_{2}$, $a_{i}.state$ is either $explore$ or $backtrack$. Note that each unsettled agent in $G_{2}$ has the same state in any round. Each agent $a_{i}$ takes $4m$ odd rounds to reach node $v^{\prime}$ and tries to move through edge $(v^{\prime},\,u^{\prime})$ as justified below.
  • if $\bm{a_{i}.state=backtrack}$: If agent $a_{i}\in G_{2}$ is with $a_{i}.state=backtrack$, then it starts a new DFS considering $u^{\prime}$ as a new root node and stores $a_{i}.skip=p$. In the next $4m$ odd rounds, it runs the DFS traversal on $G-\{e^{\prime}\}$. Therefore, it reaches node $v^{\prime}$ unless all agents of $G_{2}$ get dispersed.
  • if $\bm{a_{i}.state=explore}$: In this case, it continues its current DFS traversal by setting $a_{i}.prt\_out=(p+1)$ mod $\delta_{u^{\prime}}$. There are two cases: Case (i) the node $v^{\prime}$ is already visited node in its current DFS traversal by the round $r^{\prime}$, or Case (ii) the node $v^{\prime}$ is not visited in its current DFS traversal by the round $r^{\prime}$.
    Case (i): Let $a_{i}$ has taken $t$ odd rounds to visit node $v^{\prime}$ in the current DFS before trying to traverse through edge $e^{\prime}$. After round $r^{\prime}$, agent $a_{i}$ continues the DFS traversal via $(p+1)$ mod $\delta_{u^{\prime}}$ port. It may be possible that it does not reach node $v^{\prime}$ via the remaining part of the DFS traversal because it was unable to move by port $p$. According to our algorithm, agent $a_{i}\in G_{2}$ reaches the root node (say $w$) of the current DFS within $4m-t$ odd rounds (as to complete the current agents need $4m$ odd rounds) and, if there is some unsettled agent $a_{i}\in G_{2}$ at node $w$, it starts a new DFS by increasing $a_{i}.dfs\_label$ by 1, considering $w$ as root node and takes at most $t$ rounds to reach node $v^{\prime}$. In total, agent $a_{i}\in G_{2}$ takes $4m-t+t$=$4m$ odd rounds to reach node $v^{\prime}$.
    Case (ii): Suppose, the current DFS traversal of agent $a_{i}$ has been executed for $t^{\prime}$ odd rounds. To complete the current DFS, agent $a_{i}\in G_{2}$ takes $4m-t^{\prime}$ odd rounds more. In this case, agent $a_{i}$ reaches node $v^{\prime}$ within $4m-t^{\prime}$ odd rounds.
If at least one unsettled agent in $G_{2}$ is at node $v^{\prime}$, each agent $a_{i}\in G_{2}$ tries to move edge $(v^{\prime},\,u^{\prime})$. Whenever the movement is unsuccessful each agent $a_{i}\in G_{2}$ increases $a_{i}.count_{2}$ by 1. Since $e^{\prime}$ is missing for $16n^{2}$ consecutive odd rounds from round $r^{\prime}$, therefore for each agent $a_{i}\in G_{2}$, $a_{i}.count_{2}$ becomes $4n^{2}$ as from round $r^{\prime}$ to reach node $v^{\prime}$ it takes $4m$ odd round. In total time from $r^{\prime}$ passes by is $4m+4n^{2}<16n^{2}$ odd rounds. Therefore, $a_{i}.count_{2}$ reaches $4n^{2}$. In this case, agent $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$\&& $a_{i}.dfs\_label$ by 1, considers node $v^{\prime}$ the root and starts the new DFS from node $v^{\prime}$ as well as stores the outgoing port of edge $(v^{\prime},\,u^{\prime})$ as $a_{i}.skip$. In the next $4m\leq 4n^{2}$ odd rounds, each agent $a_{i}\in G_{2}$ completes the DFS traversal on $G-e^{\prime}$. It is because even if it reaches node $u^{\prime}$ and wants to go through edge $e^{\prime}$ via port $p$, it reaches as $a_{i}.state=explore$. According to our algorithm, agent $a_{i}$ changes the outgoing port by $(p+1)$ mode $\delta_{u}$ and continues its DFS traversal on $G-e^{\prime}$. The total odd rounds till dispersion of agents in $G_{2}$ is at most $12n^{2}$ odd rounds. Therefore, if edge $e^{\prime}$ does not appear in consecutive $16n^{2}$ rounds (we explain later why we are using $16n^{2}$ instead of $12n^{2}$), then the group $G_{2}$ is successfully dispersed in $G$.∎
If the edge $e^{\prime}$ reappears at some round before the wait time of the group $G_{1}$, then agents in $G_{1}$ move at least one round of its current DFS. Now we make a claim that guarantees the dispersion of agents in $G_{1}$.
Claim 2: For any unsettled agent $a_{i}\in G_{2}$, if $a_{i}.count_{3}$ matches with $4n^{2}$, then agents in $G_{1}$ are already dispersed.
Proof
Let $T$ be an odd round when edge $e^{\prime}$ reappears and agents in $G_{1}$ move at least one round on its current DFS traversal. Note that $T>r^{\prime}$. After round $T$ round onwards the following cases are possible.
  • Case (A) Agents in group $G_{1}$ and $G_{2}$ want to go through the same deleted edge: In this case, the group $G_{2}$ meets with $G_{1}$ at node $u_{1}$ and both want to go through the same edge $e_{1}=(u_{1},\,v_{1})$ in round $r_{1}$ (say corresponding port $p_{1}$ of $u_{1}$). There are two scenarios: (i) $G_{1}$ reaches node $u_{1}$ in round $r^{\prime\prime}<r_{1}$ but $G_{2}$ is not present, or (ii) $G_{2}$ reaches node $u_{1}$ in round $r^{\prime\prime\prime}<r_{1}$ but $G_{1}$ is not present in round $r^{\prime\prime\prime}$. In both cases, edge $e_{1}$ is deleted from round $r^{\prime\prime}$ or $r^{\prime\prime\prime}$ onwards (note that $r^{\prime\prime},r^{\prime\prime\prime}\geq T$). Within the next $4m$ odd rounds, the other group $G_{2}$ may come to node $u_{1}$ and move through edge $e_{1}$ (this is similar to Case (i) &\&& (ii) as discussed in Claim 1). These two situations change to the case that we have discussed in Claim 1. After round $r_{1}$, if the edge is deleted for $12n^{2}$ odd rounds consecutively, then $G_{2}$ gets dispersed. Due to this reason, we are giving waiting time $16n^{2}$ instead of $12n^{2}$ to $G_{1}$.
  • Case (B) Agents in $G_{2}$ want to go through deleted edge: In this case, agents in $G_{2}$ at node $u_{2}$ try to move via edge $e_{2}=(u_{2},\,v_{2})$ using port $p_{2}$. Agent $a_{i}\in G_{2}$ has parameter $a_{i}.count_{2}$ and it is increasing this parameter by 1 whenever its movement becomes unsuccessful. If its movement is successful, then $a_{i}.count_{2}=0$. If at some odd round, say $t^{\prime}$, $a_{i}.count_{2}$ reaches to $4n^{2}$ (it means $4n^{2}$ consecutive odd rounds edge $e^{\prime}$ is missing), then agent $a_{i}\in G_{2}$ restarts the new DFS and stores $a_{i}.skip=p_{2}$ corresponding edge $e^{\prime}$ through which it is trying to move. It is important to note that if edge $e_{2}$ is missing for consecutive $4n^{2}$ odd rounds, then two cases are possible which are as follows.
    • Case (a): While $G_{2}$ is waiting for $e_{2}$ to appear, then within $4n^{2}$ odd rounds unsettled agents in $G_{1}$ get stuck at node $u_{2}$ (or $v_{2}$). If each unsettled agent $a_{j}\in G_{1}$ reaches node $u_{2}$ and wants to go through edge $e_{2}$, then it is similar to Case (A). If each unsettled agent $a_{j}\in G_{1}$ gets stuck at node $v_{2}$, then it tries to move through edge $(v_{2},\,u_{2})$ and starts increasing $a_{j}.count_{1}$ parameter by 1 in each unsuccessful movement. When $a_{i}.count_{2}=4n^{2}$, then agent $a_{i}\in G_{2}$ restarts new DFS, stores $a_{i}.skip=p_{2}$ and increases parameter $a_{i}.count_{3}$ by one. If edge $e_{2}$ does not appear next odd $4m$ rounds (note that the parameter for agent $a_{j}\in G_{1}$ is less than $16n^{2}$), then agents in $G_{2}$ achieve the dispersion. It is because even if it reaches node $v_{2}$ and wants to go through edge $(v_{2},\,u_{2})$ via some port $p$, it reaches as $a_{i}.state=explore$. Agent $a_{i}$ changes the outgoing port by $(p+1)$ mode $\delta_{v_{2}}$ and continues its DFS traversal on $G-\{e_{2}\}$. Therefore, if edge $e_{2}$ does not appear in the next consecutive $4m$ odd rounds, then the group $G_{2}$ is successfully dispersed at the nodes of $G$. If edge $e_{2}$ reappears then agents in $G_{1}$ move at least one round of its current DFS.
    • Case (b): The DFS traversal of $G_{1}$ is unaffected due to the missing edge $e_{2}$. In this case, agents in $G_{1}$ get dispersed.
It is important to note that if unsettled agent $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$, then agent $a_{j}\in G_{1}$ moves at least one round of its current DFS. Therefore, whenever $a_{i}.count_{3}$ reaches $4n^{2}$, agents in $G_{2}$ can understand that agents in $G_{1}$ are dispersed as $G_{1}$ already ran its dispersion algorithm via DFS for at least $4m$ odd rounds without deviating from the path. ∎
Based on Claim 1 &\&& 2, we have the following lemma.
Lemma 1
Our algorithm solves Dispersion on TVG for $k=n+1$ agents correctly.
Proof
Due to Claim 1 &\&& 2, we can say either $G_{1}$ or $G_{2}$ is dispersed and unsettled agents in $G_{1}\,(G_{2})$ understand whether agents in $G_{2}\,(G_{1})$ are dispersed. And, agents in $G_{1}\,($or $G_{2})$ divide into two new groups. In this way, in the end, group $G_{1}$ (and $G_{2}$) contains 1 agent. Both groups start running our algorithm, due to Claim 1 &\&& 2, we can say that either agent in $G_{1}$ or $G_{2}$ is settled at some node. Therefore, our algorithm solves Dispersion on TVG for $k=n+1$ correctly. ∎
Due to Lemma 1, we can say that one group is divided into two groups only after the other group achieves dispersion. Since we have $n+1$ agents, in this way, all agents are getting dispersed except one agent which will be part of the last remaining group.
Lemma 2
The time complexity of our algorithm is $O(n^{4}\cdot\log n)$.
Proof
In Lemma 1, we have shown that the agents achieve the dispersion. If the adversary does not remove any edge in the first $4m$ odd rounds then all agents are dispersed except one agent. In odd round $r$ (where $r<4m$), if the adversary removes the edge $e$ and the group of agents wants to go through the edge $e$ according to their DFS algorithm, then they divide into two groups. We show that in the next $32n^{4}$ odd rounds, either $G_{1}$ or $G_{2}$ gets dispersed. We have proved that whenever $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$ by 1, then agents in $G_{1}$ move at least one round of its DFS or agents in $G_{1}$ get dispersed. Therefore, if agents in $G_{2}$ are not dispersed and for $a_{i}\in G_{2}$, $a_{i}.count_{3}$ reaches $4n^{2}$, then agents in $G_{2}$ can understand that agents in $G_{1}$ are dispersed. Suppose agents $G_{2}$ runs each DFS algorithm for $4m$ rounds, therefore $4(m+n^{2})\cdot 4n^{2}\leq 32n^{4}$ odd rounds are required to disperse $G_{1}$. We also have shown if $G_{1}$ waits for an edge to reappear for $16n^{2}$ rounds, then $G_{2}$ is dispersed. Therefore, after dividing into two groups, in the next $32n^{4}$ odd rounds, either $G_{1}$ or $G_{2}$ is dispersed. Therefore, either $G_{1}$ or $G_{2}$ gets dispersed in the first $32n^{4}+4m$ odd rounds. Including even rounds, we can say that after dividing into two groups, in the next $64n^{4}+8m$ rounds, either $G_{1}$ or $G_{2}$ is dispersed.
Using Lemma 1, if $G_{1}$ (or $G_{2}$) understands that $G_{2}$ (or $G_{1}$) is dispersed, then $G_{1}$ (or $G_{2}$) divides into two new groups, and both groups contain half agents of $G_{1}$ (or $G_{2}$). These two groups repeat the same procedure. The number of such group divisions is $O(\log n)$, and one of two groups is dispersed in $64n^{4}+8m$ rounds. Therefore, our algorithm takes $O(n^{4}\cdot\log n)$ rounds to solve Dispersion on TVG for $k=n+1$. ∎
The following lemma is w.r.t. memory lower bound. The proof of this lemma is motivated from [2]. In [2], they have provided the memory lower bound to solve dispersion on a static graph when $k=n$. We extend their idea for the memory lower bound when $k>n$.
Lemma 3
Assuming all agents are given the same amount of memory, agents require $\Omega(\log n)$ bits of memory each for any deterministic algorithm to solve Dispersion on TVG when $k\geq n+1$.
Proof
Suppose all agents have $o(\log n)$ bits of memory. Each agent’s state space is $2^{o(\log n)}=n^{o(1)}$. Since there are at least $n+1$ agents, by pigeonhole principle, there exist at least $k-n+2$ agents with the same state space. Let agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ have the same state space. Let us suppose that all agents are initially co-located at some node. Since all agents run the same deterministic algorithm, and agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are co-located initially, they will run the same moves. It means that agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ remain stay together. Therefore, by the end of the algorithm, agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are at the same node. Solving Dispersion on TVG requires a configuration where there is at least one agent at every node. We will never solve Dispersion on TVG because there is no way for agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ to settle down on at least two different nodes. Therefore, all $k$ agents are at most $n-1$ nodes by the end of the algorithm. Thus, to solve Dispersion on TVG, agents need $\Omega(\log n)$ memory each. ∎
Lemma 4
In our algorithm, agents require $\Theta(\log n)$ memory.
Proof
For agent $a_{i}$, $a_{i}.grp\_label$ is nothing but a binary string. Whenever we divide the group, the $a_{i}.grp\_label$ is appended by 0 or 1. The number of group divisions is $O(\log n)$. Therefore, the length of $a_{i}.grp\_label$ is $O(\log n)$, which agent $a_{i}$ can store in its $O(\log n)$ memory. According to Lemma 2, all agents achieve dispersion in $O(n^{4}\cdot\log n)$ rounds. Therefore, $a_{i}.dfs\_label$ can not be more than $O(n^{4}\cdot\log n)$ which agent $a_{i}$ can store in its $O(\log n)$ memory. Apart from these two parameters, all unsettled agents $a_{i}$ either store boolean variables or port information. Such parameters are constant as per our algorithm. To store each parameter, we need $O(\log n)$ storage. Therefore, each unsettled agent $a_{i}$ needs $O(\log n)$ memory. All settled agents maintain two parameters as per our algorithm $a_{i}^{u}(G_{1})$ and $a_{i}^{u}(G_{2})$. To store each parameter, settled agents need $O(\log n)$ memory. Therefore, each settled agent $a_{i}$ needs $O(\log n)$ memory.
Due to Lemma 3, the memory lower bound for the dispersion of $k\geq n+1$ agents is $\Omega(\log n)$, and in our algorithm agents use $O(\log n)$ memory each. Therefore, the agents in our algorithm require $\Theta(\log n)$ memory due to $k=n+1$. ∎
Now we are ready to provide our main theorem  [^1] Our algorithm can be extended to solve Dispersion on TVG when $k>n$. Since agents are co-located they can consider $k$ as an upper bound of $n$. In our algorithm in place of $n$, if we use $k$, our algorithm solves Dispersion on TVG with $k>n+1$ co-located agents and the time complexity is $O(k^{4}\cdot\log n)$..
Theorem 3.1
Our algorithm solves Dispersion on TVG for $k=n+1$ in $\mathcal{G}$ in $O(n^{4}\cdot\log n)$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The proof follows from Lemma 1, 2, and 4. ∎
Observation 1
All agents know the value of $n$, and $m\leq n^{2}$. If all agents want to achieve the termination, can calculate the value of $72n^{4}\cdot\lceil\log n\rceil$ and terminate after $72n^{4}\cdot\lceil\log n\rceil$ rounds.
4 Conclusion
In this work, we have studied the Dispersion on TVG on $\mathcal{G}$. We found that with $k=n$ agents starting from a rooted configuration, 1-hop visibility and global communication are necessary to solve the Dispersion on TVG. Additionally, we have shown that if $k\geq n+1$, agents do not require 1-hop visibility and global communication to solve the Dispersion on TVG from any rooted initial configuration. One can try to improve the time complexity for the case $k\geq n+1$. It would also be interesting to explore the Dispersion on TVG for $k\geq n+1$ when agents start from arbitrary initial configuration. Another interesting question that remains open is whether global communication and 1-hop visibility are necessary to solve dispersion for $k<n$ agents.