### v2

# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

Abstract
Besides being studied over static graphs heavily, the dispersion problem is also studied on dynamic graphs with $n$ nodes where at each discrete time step the graph is a connected sub-graph of the complete graph $K_{n}$. An optimal algorithm is provided assuming global communication and 1-hop visibility of the agents. How this problem pans out on Time-Varying Graphs (TVG) is kept as an open question in the literature. In this work, we study this problem on TVG considering $k\geq 1$ agents where at each discrete time step the adversary can remove at most one edge keeping the underlying graph connected. We have the following main results considering all agents start from a rooted initial configuration. Global communication and 1-hop visibility are must to solve dispersion for $pn$ ($p\geq 1$) co-located agents on a TVG even if agents have unlimited memory and knowledge of $n$. We provide an algorithm that disperses $n+1$ agents on TVG by dropping both the assumptions of global communication and 1-hop visibility using $O(\log n)$ memory per agent. We extend this algorithm to solve dispersion with $pn+1$ agents with the same model assumptions.
Keywords:
Dispersion Mobile agents Distributed algorithms Time-Varying Graph Deterministic algorithms
1 Introduction
The dispersion problem, introduced in [3] on static graphs, involves the coordination of $k\leq n$ agents on the graph of size $n$ to reach a configuration with the constraint that at most one agent can be present at any node. This problem generally applies to real-world scenarios in which agents must coordinate and share resources at various places. The goal is to minimize the total cost of solving the problem while ensuring that the cost of agents moving around on the graph is much lower than the cost of having more than one agent share a resource. The dispersion problem is also connected to three other problems: (i) scattering or uniform deployment of agents on a graph [8], (ii) exploration of agents on a graph [9], and (iii) load balancing on a graph [6]. The scattering problem involves coordinating and moving $k\leq n$ agents on a given graph with $n$ nodes so that they are uniformly distributed over the graph. When $k=n$, the scattering problem is equivalent to the dispersion problem. The problem of collaborative exploration requires agents to visit all the nodes of a given graph while coordinating with one another. This problem is related to dispersion since if $n$ agents need to disperse, they must explore the graph. Load balancing on graphs is a problem where resources need to be evenly distributed as much as possible among nodes on a graph in the least possible number of rounds. Typically, nodes control the allocation of resources, and there are constraints on how much load can pass through an edge in a round. This problem is similar to dispersion if we consider agents as resources.
In [21], Molla et al. examine the dispersion problem for the first time for any value of $k\geq 1$. In this work, they have given the following definition of the dispersion.
Definition 1
Let $k$ agents be present at $n$ node graph $G$. Each node should have at most $\lceil k/n\rceil$ agents in the final configuration.
There are several works on dispersion [3, 13, 14, 22, 17, 23, 21], and most of them deal with static graphs. The addition of dynamism to a network makes the dispersion problem more challenging as well as practical. Agents need to complete their tasks in environments that are also changing. In this work, we aim to solve the dispersion problem for any $k$ on dynamic graphs. Below, we discuss the model assumption and the problem definition.
1.1 Model and the problem
Time-varying graph model: The system is represented as a time-varying graph (TVG), denoted by $\mathcal{G}=(V,E,\mathbb{T},\rho)$, where, $V$ represents the set of nodes, $E$ represents a set of edges, $\mathbb{T}$ represents the temporal domain, which is defined to be $\mathbb{Z}^{+}$ as in this model we consider discrete time steps, and $\rho:E\times\mathbb{T}\rightarrow\{0,1\}$ tells whether a particular edge is available at a given time. The graph $G=(V,\,E)$ is the underlying static graph of the dynamic graph $\mathcal{G}$, also termed as a footprint of $\mathcal{G}$. We denote $\delta_{v}$ as the degree of node $v$ in the footprint $G$, and $\Delta$ as the maximum degree in the footprint $G$. The graph $G$ is undirected and unweighted. The nodes of $G$ are anonymous (i.e., they have no IDs). The graph $G$ is port-labelled, i.e., each incident edge of a node $v$ is assigned a locally unique port number in $\{0,1,2,3,\cdots,\delta_{v}-1\}$. Specifically, each undirected edge $(u,\,v)$ has two labels, denoted as its port numbers at $u$ and $v$. Port numbering is local, i.e., there is no relationship between the port numbers at $u$ and $v$. We consider there is an adversary that can remove edges from the footprint $G$, keeping $\mathcal{G}$ connected at each round.
Agent: We consider $k$ agents to be present initially at nodes of the graph $G$. Each agent has a unique identifier assigned from the range $[1,\,n^{c}]$, where $c$ is a constant. Each agent knows its ID. Agents are not aware of the values of $n$, $\Delta$, or $c$ unless these values are specifically mentioned. The agents are equipped with $O(\log n)$ memory. An agent residing at a node (say $v$) in round $t$ knows $\delta_{v}$ in the footprint $G$ and all associated ports corresponding to node $v$ in $G$; however, the agent does not understand if any incident edge of node $v$ is missing in $\mathcal{G}$ in round $t$. To be more precise, agent $a$ currently positioned at node $v$ at round $t$ does not know the value of $\rho(e_{v},t)$, where $e_{v}$ is an edge incident at node $v$. Such a model has been considered in [11, 12].
Communication model: There are two communication models in a distributed setting: local and global. In the local communication model [3, 13], an agent located at a particular graph node can communicate with other agents present at the same node. In contrast, the global communication model [10, 16] allows an agent at a graph node to communicate with any other agent in the graph, even if they are located at different nodes. In nature, this type of communication happens through the links of the graph. We use the local communication model in our algorithm.
Visibility model: In this work, we use two types of visibility models: zero-hop visibility and $1$-hop visibility. In the zero-hop visibility model [3, 14, 15], an agent located at a particular node (say $v\in\mathcal{G}$) can see agents present at node $v$ in round $r$, and port numbers associated with node $v$. In $1$-hop visibility [1, 20], an agent (say $a_{i}$) located at a particular node (say $v\in\mathcal{G}$) can see any node (along with agent(s) present at that node, if any) that is the neighbours of $v$ in $\mathcal{G}$. If the adversary deletes an edge associated with a node $u$ with port $p$ at the start of a round $r$, the agent(s) present there cannot understand that an edge has been deleted in the zero-hop visibility model. However, if the agent tries to move through that edge in this round, it cannot make the move. In the 1-hop visibility model, if such an edge is deleted in a round, then the agents present there can see all the neighbours of $v$ that are connected to $v$ via the edges with the port number $\neq p$ at $v$ because according to the 1-hop visibility model, the agents can see the sub-graph induced by the nodes that are 1-hop away in this round. Therefore, agents can also recognize that no node is connected through port $p$, and therefore, agents can understand that the edge corresponding to port $p$ at $v$ is deleted at the start of this round.
The algorithm runs in synchronous rounds. In each round $t$, an agent $a_{i}$ performs one Communicate-Compute-Move (CCM) cycle as follows:
  • Communicate: Agent $a_{i}$ at node $v_{i}$ can communicate with other agents $a_{j}$ present at the same node $v_{i}$ or present at any different node $v_{j}$, depending on the communication model used. The agent also understands whether it had a successful or an unsuccessful move in the last round.
  • Compute: Based on the information the agent has, the agent computes the port through which it will move or decides not to move at all.
  • Move: Agent moves via the computed port or stays at its current node.
Now, we are ready to provide our problem definition.
Definition 2
$k$-Dispersion on TVG: Let $\mathcal{G}$ be a TVG. Let $k$ $(\geq 1)$ agents be at the nodes of the footprint $G$. At each round, the adversary can remove at most one edge from $G$, keeping $\mathcal{G}$ connected. The agents need to reposition themselves such that each node contains at most $\lceil k/n\rceil$ agents.
In our problem definition, we consider the adversary can delete at most one edge from footprint $G$ at each round and this restriction is already considered in several existing works [4, 5, 12].
ID Range IC Dynamism $\bm{k}$ Necessary Assumptions Sufficient Assumptions Algorithm
[16] [1, k] Scattered Arbitrary edge addition or deletion maintaining the graph connected at each round $k\leq n$ 1-hop visibility, global comm. 1-hop visibility, global comm. $\Theta(k)$ rounds, $\Theta(\log k)$ memory per agent
This work $[1,\,n^{c}]$, $c\in\mathbb{N}$ Co-located 1-TVG $k=n+1$ 0-hop visibility, local comm. $O(n^{4}\cdot\log n)$ rounds, $\Theta(\log n)$ memory per agent
$k=pn+1$, $p>1$ 0-hop visibility, local comm., knowledge of $n$ $O(n^{4}\cdot\max(\log n,\,\log p))$ rounds, $\Theta(\log n)$ memory per agent
Co-located 1-TVG $k=pn$, $p\geq 1$ 1-hop visibility, global comm.
Scattered $f$-TVG $k\geq 1$ 1-hop visibility, global comm. $O(k)$ rounds, $\Theta(\log n)$ memory per agent (An extension of [16])
Table 1: This table shows the results of dispersion on different dynamic graph models. In this table, IC represents the initial configuration, $1$-TVG means the adversary can delete at most one edge per round from footprint $G$ while maintaining $\mathcal{G}$ connected, and $f$-TVG means the adversary can delete at most $f$ edges per round from footprint $G$ while maintaining $\mathcal{G}$ connected.
1.2 Related work
The problem of dispersion for $k$ $(\leq n)$ agents has been extensively studied on static graphs [3, 13, 14, 22, 17, 23]. Here we focus on the literature of dispersion on dynamic graphs only. Recently, researchers in the field of distributed computing started studying dynamic graphs, where the topological changes are not sporadic or anomalous but rather inherent to the nature of the network. Such a dynamic graph model was developed by Kuhn et al. [18] in 2010. Dispersion on the dynamic graphs for $k\leq n$ agents is studied by Agarwalla et al. [1] and Kshemkalyani et al. [16]. In [1], the authors consider the following dynamic graph model on ring $\mathcal{R}$. In [1], the adversary can delete at most one edge from $\mathcal{R}$ at each round. Based on this dynamic graph model, they develop several deterministic algorithms for agents to achieve dispersion on a ring. In [16], the authors use a weaker dynamic graph model than [1]. In [1], the adversary can delete at most one edge per round from the ring, but in [16], the adversary can add or remove edges, keeping the dynamic graph connected at each round. The authors show that it is impossible to solve dispersion on the dynamic graph in the local communication model, even if $1$-hop visibility is available to the agents and each agent has unlimited memory. Also, it is impossible to solve dispersion on the dynamic graph in the global communication model without $1$-hop visibility, even with unlimited memory at each agent. They provided an asymptotically optimal $\Theta(k)$ rounds algorithm in the global communication model with $1$-hop visibility.
In [2], Casteigts et al. defined the new dynamic graph paradigm, which is known as TVG. In this model, the exploration [11], gathering [19], and black-hole search problem [7, 4, 5, 12] have been studied. The basic difference between the dynamic graph model used in [16] and TVG is as follows. In [16], the footprint is a clique and for TVG’s the footprint can be graphs with less number of edges as well. Moreover, after edge deletion/addition, port numbers are adjusted in the model of [16], which is not the case in the TVG model. This allows us to understand missing edges in the presence of 1-hop visibility. That’s why TVG exhibits a stronger graph model than the dynamic graph model used in [16]. The authors [16] ask the following question on the behaviour of the dispersion problem on TVG. Since the graph model of TVG is stronger than the dynamic model used in [16], their impossibility results do not hold here on TVGs.
In this work, we study our $k$-Dispersion on TVG. In the preliminary version of this work, we demonstrated that 1-hop visibility and global communication are necessary for solving $n$-Dispersion on TVG, and we provided a high-level approach for addressing $(n+1)$-Dispersion on TVG. Here, we extend those results and provide our algorithms in detail.
1.3 Our contribution
In this paper, we provide the following results.
  • Considering 1-hop visibility but without global communication, it is impossible to solve $k$-Dispersion on TVG when $k=pn$, $p\in\mathbb{N}$. This impossibility is valid even if the agents start from a rooted configuration, have knowledge of $n$ and are equipped with infinite memory (refer Section 2).
  • Considering global communication but without 1-hop visibility, it is impossible to solve $k$-Dispersion on TVG when $k=pn$, $p\in\mathbb{N}$. This impossibility is valid even if the agents start from a rooted configuration, have knowledge of $n$ and are equipped with infinite memory (refer Section 2).
  • We provide $\Omega(\log n)$ memory lower bound per agent to solve $k$-Dispersion on TVG (refer to Section 2), when $k>n$.
  • We present an algorithm that solves $(n+1)$-Dispersion on TVG when $n+1$ agents are initially co-located. This algorithm takes $O(n^{4}\cdot\log n)$ rounds and requires $O(\log n)$ memory per agent, considering local communication and zero visibility. This algorithm is memory-optimal (refer Section 3). If agents know $n$ and, $pn+1$ agents are co-located, then the same algorithm can be extended to solve $(pn+1)$-Dispersion considering local communication and zero visibility, where $p\in\mathbb{N}$ (refer to Section 4.1).
  • To solve $k$-Dispersion on TVG, we can use the existing algorithm of [16] when agents are equipped with 1-hop visibility and global communication. Further, the same algorithm of [16] with minor modifications works even when the adversary can delete any number of edges while keeping $\mathcal{G}$ connected in each round and $k(\geq 1)$ agents start from arbitrary initial configuration (refer to Section 4.2).
Figure 1: (a) If the adversary does not delete any edge at the beginning of round $r$, then this shows the computation of the agents. (b) If the adversary deletes $(u,\,u_{2})$ at the beginning of round $r$, then the view of agents at node $u_{1}$ does not change, but the view of agents at node $u$ and $u_{1}$ changes due to 1-hop visibility. In this case, agents at node $u$ can change $x_{1}$ to $x_{11}$, and agents at node $u_{2}$ can change $y_{3}$ to $y_{33}$ but agents at node $u_{1}$ do not change their decision because they can not understand deletion of edge $(u,\,u_{2})$ due to unavailability of global communication. Based on the cases mentioned in Theorem 2.1, the adversary decides which edge to delete at the start of round $r$. This outlines the high-level idea behind our proof Theorem 2.1.
2 Impossibility Results and Lower Bound
In this section, we show that it is impossible to solve $k$-Dispersion on TVG deterministically in our model if either global communication or 1-hop visibility is omitted, even if agents start from a rooted configuration. These impossibility results also hold if agents have infinite memory and knowledge of $n$. Recall that according to our model, the agents at a node $v$ cannot know whether an incident edge to node $v$ is missing or not unless agents are equipped with $1$-hop visibility.
Theorem 2.1
(Impossibility without global communication) It is impossible to solve $k$-Dispersion on TVG for any $n\geq 6$ with the agents having $1$-hop visibility but no global communication even if agents are initially co-located. This impossibility result also holds if agents have infinite memory and knowledge of $n$.
Proof
We prove this impossibility on TVG ring $\mathcal{R}$. Let $k=pn$ many agents be co-located at some node $v$ of ring $\mathcal{R}$, where $p$ is any positive integer. If the agents have 1-hop visibility, then as per our visibility model, the agent can sense the missing edge corresponding to some port $p$. There is no global communication. Therefore, the movement of the agents depends on 1-hop visibility and local communication. Since the algorithm is deterministic, the adversary can precompute the agents computation and accordingly remove an edge from $G$. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movements of agents do not lead to dispersion, the adversary does not need to do anything. If the movements of agents lead to dispersion, we discuss how the adversary intervenes. Assume that at the end of round $r$, the agents successfully achieve dispersion. It means that the dispersion is not achieved at the beginning of round $r$, and after the CCM cycle at round $r$, the agents are in the dispersed configuration. Since $k=pn$, each node of $\mathcal{R}$ should have $p$ many agents in the dispersed configuration. At the beginning of round $r$, there is a node $u$ which has at most $p-1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in ring $\mathcal{R}$. Let $z$ ($\leq p-1$) agents be at node $u$. If the adversary does not delete any edge at the beginning of round $r$, then the following is the computation of the agents according to $\mathcal{A}$: let $x_{1}$ agents move from node $u$ to $u_{1}$, $y_{1}$ agents move from node $u$ to $u_{2}$, $x_{2}$ agents move from node $u_{1}$ to $u$, and $y_{2}$ agents move from node $u_{2}$ to $u$. And accordingly, at the end of round $r$, there are $z+(x_{2}-x_{1})+(y_{2}-y_{1})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$ and accordingly deletes an edge. A high-level idea behind the impossibility proof is given in Figure 1. We divide the cases as follows.
  • Case $\bm{x_{2}=0}$ or $\bm{y_{2}=0}$: W.l.o.g. let $x_{2}=0$. In this case, the adversary deletes the edge $(u,\,u_{2})$ at the beginning of round $r$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change. Also, since global communication is not there, agents at node $u_{1}$ can not get the information about the missing edge $(u,\,u_{2})$ and hence the movement of agents at node $u_{1}$ will remain the same in round $r$. But the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Therefore, at the end of round $r$, there are $z-x_{1}^{\prime}<p$ agents at the end of round $r$, where $x_{1}^{\prime}$ is the number of agents moving from node $u$ to $u_{1}$ due to the computation at $u$ based on the removed edge. Therefore, at the end of round $r$, the dispersion configuration is not achieved.
  • Case $\bm{x_{1}=0}$ or $\bm{y_{1}=0}$: W.l.o.g. let $x_{1}=0$. At the end of round $r$, there are $z+x_{2}+(y_{2}-y_{1})=p$ agents at node $u$. It is important to note that $x_{2}>0$ and $y_{2}>0$. Based on the values of $y_{1},\,y_{2}$, we have the following cases.
    $\rightarrow$ Sub-case $\bm{y_{1}=y_{2}:}$ The adversary deletes the edge $(u,\,u_{1})$. Therefore, the 1-hop view of agents at node $u_{2}$ does not change, but the 1-hop view of agents at node $u$ and $u_{1}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{2}$ will remain the same in round $r$. In this case, from node $u$, more than $y_{1}$ agents or less than $y_{1}$ agents can not move to node $u_{2}$; otherwise, node $u_{2}$ contains either less than $p$ agents or greater than $p$ agents. If exactly $y_{1}$ agents move to node $u_{2}$, then the number of agents at node $u$ is $z$, and $z$ is less than $p$. Therefore, the dispersion configuration is not achieved at the end of round $r$.
    $\rightarrow$ Sub-case $\bm{y_{1}\neq y_{2}:}$ The adversary deletes the edge $(u,\,u_{2})$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change, but the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{1}$ will remain the same in round $r$. In this case, if agents move from node $u$ to $u_{1}$, then there are at least $p+1$ agents at node $u_{1}$ due to the view of node $u_{1}$. Therefore, the number of agents at node $u$ is $z+x_{2}$. If $z+x_{2}=p$, then $z+x_{2}=z+x_{2}+(y_{2}-y_{1})\implies y_{1}=y_{2}$. This gives a contradiction. Therefore, node $u$ does not contain $p$ agents at the end of round $r$. Therefore, the dispersion is not achieved in this case as well at the end of round $r$.
  • Case $\bm{x_{1},\,x_{2},\,y_{1},\,y_{2}>0}$: It is important to note that $(x_{2}-x_{1})\leq 0$ and $(y_{2}-y_{1})\leq 0$ are not possible. Therefore, either $(x_{2}-x_{1})>0$ or $(y_{2}-y_{1})>0$. W.l.o.g., let $(y_{2}-y_{1})>0$. In this case, the adversary deletes the edge $(u,\,u_{2})$ at the beginning of round $r$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change, but the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{1}$ will remain the same in round $r$. In this case, if more than $x_{1}$ or less than $x_{1}$ agents move from $u$ to $u_{1}$, then at node $u_{1}$, the number of agents is not $p$ due to the fact the view of node $u_{1}$. Therefore, the number of agents at node $u$ is $z+(x_{2}-x_{1})$. If $z+(x_{2}-x_{1})=p$, then $z+(x_{2}-x_{1})=z+(x_{2}-x_{1})+(y_{2}-y_{1})\implies(y_{2}-y_{1})=0$. This gives a contradiction. Therefore, at the end of round $r$, the number of agents at node $u$ is not $p$. Hence, the dispersion is not achieved at the end of round $r$.
This completes the proof. ∎
Figure 2: (a) If the adversary does not delete any edge at the beginning of round $r$, then this shows the computation of the agents. (b) If the adversary deletes the edge $(u,u_{2})$ at the beginning of round $r$, then the views of the agents at nodes $u_{1}$, $u_{2}$, and $u_{3}$ remain unchanged because they cannot detect the missing edge. In this situation, the agents at nodes $u$, $u_{1}$ and $u_{2}$ will attempt to send the same number of agents, but their movement will not be successful in round $r$. This fact makes the proof of Theorem 2.2 easy. Based on the cases mentioned in Theorem 2.2, the adversary decides which edge to delete at the start of round $r$. This outlines the high-level idea behind the proof of Theorem 2.2.
Theorem 2.2
(Impossibility without 1-hop visibility) It is impossible to solve $k$-Dispersion on TVG for any $n\geq 6$ considering global communication but without $1$-hop visibility even if agents are initially co-located. This impossibility result also holds if agents have infinite memory and knowledge of $n$.
Proof
We prove this impossibility on TVG ring $\mathcal{R}$. Let $k=pn$ many agents be co-located at some node $v$ of ring $\mathcal{R}$, where $p$ is any positive integer. If the agents have zero-hop visibility, then the choice of movement of each agent depends only on the information present at the current node and the information attained via global communication. Since the algorithm is deterministic, the adversary can remove an edge from $G$ based on the movement of the agents. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movement of agents does not lead to dispersion, the adversary does not need to do anything. If the movement of agents leads to dispersion, the adversary has to intervene. Assume that at the end of round $r$, the agents successfully achieve dispersion. It means that the dispersion is not achieved at the beginning of round $r$, and after the CCM cycle at round $r$, the agents are in the dispersed configuration. Since $k=pn$, each node of $\mathcal{R}$ should have $p$ many agents in the dispersed configuration. At the beginning of round $r$, there is a node $u$ which has at least $p+1$ many agents, and there is a node $v$ which has at most $p-1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in ring $\mathcal{R}$. Based on the agent’s movement from node $u$, we give the strategy for the adversary at the beginning of round $r$ such that the dispersed configuration is not achieved at the end of round $r$. A high-level idea behind the impossibility proof is given in Figure 2. We divide the cases as follows.
  • Case 1. According to $\mathcal{A}$, if all the agents at node $u$ stay after $CCM$, then the adversary does not delete any edge at the beginning of round $r$. Since the agents of node $u$ do not move, therefore, at node $u$, there are at least $p+1$ agents at the end of round $r$. Hence, the dispersed configuration is not achieved at the end of round $r$.
  • Case 2. According to $\mathcal{A}$, at least one agent at node $u$ move to $u_{1}$ or $u_{2}$ after $CCM$ at round $r$. W.l.o.g., at least one agent at node $u$ moves to node $u_{1}$, and no agents move from $u$ to $u_{2}$. Let $x$ agents move to node $u_{1}$. The adversary deletes edge $(u,\,u_{1})$ at the beginning of round $r$. Since agents have 0-hop visibility, they can not understand the deletion of the edge at round $r$. Since edge $(u,u_{1})$ has been deleted, the movement of agents will be unsuccessful. In this case, at the end of round $r$, at node $u$, there will be at least $p+1$ agents. Hence, the dispersed configuration is not achieved at the end of round $r$.
  • Case 3. Let $z$ agents be at node $u$ at the beginning of round $r$, where $z>p$. According to $\mathcal{A}$ at round $r$, at least one agent at node $u$ moves to $u_{1}$, and at least one agent at node $u$ moves to node $u_{2}$. Let $x_{1}$ agents from node $u$ move to node $u_{1}$, and $y_{1}$ agents from node $u$ move to node $u_{2}$. Note that $x_{1},\,y_{1}\geq 1$. Let $x_{2}$ agents from node $u_{1}$ move to node $u$, and $y_{2}$ agents from node $u_{2}$ move to node $u$, where $x_{2},\,y_{2}\geq 0$. If the adversary does not delete any edge, then at the end of round $r$, there are $(z-x_{1}-y_{1}+x_{2}+y_{2})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$ and accordingly deletes an edge. In this case, we divide this case into two sub-cases.
    • Sub-case $\bm{x_{1}=x_{2}}$ and $\bm{y_{1}=y_{2}:}$ At the end of round $r$, there are $z$ agents at node $u$, and $z>p$. Therefore, the dispersion configuration is not achieved at the end of round $r$.
    • Sub-case $\bm{x_{1}\neq x_{2}}$ or $\bm{y_{1}\neq y_{2}:}$ W.l.o.g., let $x_{1}\neq x_{2}$. In this case, the adversary deletes the edge $(u,\,u_{1})$ at the beginning of round $r$. Therefore, the movement of agents from node $u$ to $u_{1}$ and node $u_{1}$ to $u$ becomes unsuccessful. Since agents are unaware of the deletion of edge, the movement of agents from node $u$ to $u_{2}$ and node $u_{2}$ to $u$ becomes successful. Therefore, the number of agents at node $u$ is $z-y_{1}+y_{2}$. To achieve the dispersed configuration at the end of round $r$, the value of $z-y_{1}+y_{2}$ must be $p$. Therefore, $z-y_{1}+y_{2}=z-x_{1}-y_{1}+x_{2}+y_{2}\implies x_{1}=x_{2}$, and it is a contradiction due to $x_{1}\neq x_{2}$. Therefore, the value of $z-y_{1}+y_{2}$ is either less than $p$ or greater than $p$. Hence, the dispersion configuration is not achieved at the end of round $r$.
This completes the proof. ∎
Remark 1
It is important to note that the proof of both impossibility results is valid when $k=pn$, where $p$ is some constant, i.e., agents need 1-hop visibility and global communication to solve $pn$-Dispersion on TVG. In Section 3, we provide an algorithm to solve $(n+1)$-Dispersion on TVG with local communication and zero-hop visibility, and extend this idea in Section 4.1 to solve $(pn+1)$-Dispersion on TVG. The aforementioned proofs do not hold when $k\in[pn+2,\,(p+1)(n-1)]$, where $p\in\mathbb{N}$. Therefore, it is an open question whether 1-hop visibility and global communication are necessary to solve $k$-Dispersion on TVG for $k\in[pn+2,\,(p+1)(n-1)]$, where $p\in\mathbb{N}$. Also, it remains open to come up with the necessary conditions to solve $k$-Dispersion on TVG for $k<n$ given the algorithm of [16] that works with global communication and 1-hop visibility is a valid solution.
Remark 2
Note that both proofs are valid in the dynamic graph model mentioned in [16] since TVG is a stronger graph model. Therefore, agents need 1-hop visibility and global communication to solve the dispersion of $pn$ agents in the dynamic graph model mentioned in [16] as well.
Now, we provide the memory lower bound result to solve $k$-Dispersion on TVG. The proof of this lemma is motivated from [3]. In [3], they have provided the memory lower bound $\Omega(\log k)$ to solve dispersion on a static graph when $k\leq n$. We extend their idea for the memory lower bound when $k>n$.
Lemma 1
Assuming all agents are given the same amount of memory, agents require $\Omega(\log n)$ bits of memory each for any deterministic algorithm to solve $k$-Dispersion on TVG when $k\geq n+1$, and this proof is valid even if the adversary does not delete any edge at any round from footprint $G$.
Proof
Suppose all agents have $o(\log n)$ bits of memory. Each agent’s state space is $2^{o(\log n)}=n^{o(1)}$. Since there are at least $n+1$ agents, by pigeonhole principle, there exist at least $k-n+2$ agents with the same state space. Let agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ have the same state space. Let us suppose that all agents are initially co-located at some node. Since all agents run the same deterministic algorithm, and agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are co-located initially, they will run the same moves. It means that agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ remain stay together. Therefore, by the end of the algorithm, agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are at the same node. Solving dispersion requires a configuration where there is at least one agent at every node. We will never solve dispersion because there is no way for agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ to settle down on at least two different nodes. Therefore, all $k$ agents are at most $n-1$ nodes by the end of the algorithm. Thus, to solve the dispersion problem, agents need $\Omega(\log n)$ memory each. This proof is for static graphs, and $TVG$ is a weaker model than static. Therefore, the same proof is also valid for TVGs. ∎
3 The Dispersion on General Graphs
In this section, we introduce an algorithm that solves the $(n+1)$-Dispersion on TVG when $n+1$ agents are co-located at some node in a general graph, and agents do not have global communication and 1-hop visibility, i.e., we assume that agents are equipped with zero-hop visibility and local communication. We use the idea of depth-first search (DFS) traversal by mobile agents that also solves the dispersion on static graphs [3]. For the sake of completeness, let us recall the idea of DFS traversal by a group of agents. At each node $v$, a group of agents chooses an unexplored port $p$ and traverses through it: (i) if it reaches a node $u$ which is an empty node (agents can understand this by the fact that there is no agent settled at $u$), then one agent settles at $u$, and stores port $p^{\prime}$ (through which it entered node $u$) as its parent port. Now, the unsettled agents try to move through port $(p^{\prime}+1)$ mod $\delta_{u}$ in the explore state, (ii) else, if the group of agents reach a node which already has a settled agent, the agent returns to node $v$ in the backtrack state. If the unsettled agents reach node $v$ in the backtrack state, then the agents compute $q=(p+1)$ mod $\delta_{v}$. If port $q$ is unexplored, it moves through port $q$ in the explore state. Else, if port $q$ equals the parent port of $v$, then the agents move through port $q$ in the backtrack state. This algorithm takes $4m$ rounds to run on any graph, where $m$ is the number of edges in that graph. The information required for the group of agents at each node to run this algorithm can be stored by the settled agent that is already present at that node.
Recall that, according to our model, an agent does not know if any associated edge w.r.t. its current position is missing or not. This makes the adversary powerful as it can precompute the agents’ decision at a node and accordingly can delete an associated edge at the start of that round. If the group of unsettled agents decides to move through port $p$ according to our algorithm in round $r$, the adversary can delete the edge corresponding to port $p$ in round $r$. Thus preventing the group of agents from moving in round $r$.
High-level idea: No group of agents moves indefinitely through which its movement was unsuccessful. Initially, there is only one group as the configuration is rooted. Whenever the group has an unsuccessful move, the agents divide into two groups, namely $G_{1}$ and $G_{2}$. These two groups run their dispersion algorithm via DFS separately. A missing edge can make the movement unsuccessful for both groups. On unsuccessful movement, $G_{1}$ never deviates from its DFS traversal path, whereas $G_{2}$ deviates after some finite waiting period. This keeps the algorithm running, and either $G_{2}$ gets dispersed, or $G_{1}$ can move one hop according to its DFS traversal. If $G_{2}$ gets dispersed, $G_{1}$ can understand it within some finite rounds and gets divided into two groups further. Else, if $G_{1}$ gets dispersed by moving one hop again and again, $G_{2}$ can understand it within some finite rounds and gets divided into two groups further. We ensure that none of $G_{1}$ (or $G_{2}$) divides further unless $G_{2}$ (or $G_{1}$) is already dispersed. This keeps the total number of groups bounded by 2. As the settled agents keep the information required for the DFS traversal of all the groups, it helps the agents store the information using $O(\log n)$ memory. As we have different algorithms for the groups and different criteria for the groups to be divided, while dividing a group into two, each new group is given a unique label so that $G_{1}\,(G_{2})$ can execute their respective algorithms.
3.1 Algorithm for general graph
In this section, we provide a detailed description of the algorithm for dispersion on a time-varying graph $\mathcal{G}$ with at most one dynamic edge. In our algorithm, unsettled agents may get divided into two groups, we call them $Group_{1}$, and $Group_{2}$ and denote them by $G_{1}$, and $G_{2}$ respectively. At any particular round, there can be no more than two groups. However, over time, if all the unsettled agents of one group get settled, the other group can be divided into two groups again. So, if we look throughout the run of the algorithm, there may exist several groups. To identify these groups, we provide unique labelling to each group whenever a group gets generated and store it in a variable. Now we provide the parameters maintained by each agent $a_{i}$.
  • $\bm{a_{i}.settled}$: It is a binary variable and indicates whether $a_{i}$ is settled at the current node or not. If $a_{i}.settled=0$ then it is not settled whereas $a_{i}.settled=1$ denotes that agent $a_{i}$ is settled. Initially $a_{i}.settled=0$.
  • $\bm{a_{i}.prt\_in}$: It stores the port through which an agent $a_{i}$ enters into the current node. Initially, $a_{i}.prt\_in=-1$.
  • $\bm{a_{i}.prt\_out}$: It stores the port through which an agent $a_{i}$ exits from the current node. Initially, $a_{i}.prt\_out=-1$.
  • $\bm{a_{i}.state}$: It denotes the state of an agent $a_{i}$ that can be either explore or backtrack. Initially $a_{i}.state=explore$.
  • $\bm{a_{i}.dfs\_label}$: It stores the number of new DFS traversals started by agent $a_{i}$. Initially $a_{i}.dfs\_label=1$.
  • $\bm{a_{i}.success}$: When an agent $a_{i}$ attempts to move from a node $u$ to $v$ via the edge $(u,\,v)$, in the round $t$, it is either successful or unsuccessful. If the agent reaches the node $v$ in the round $t+1$, it is a successful attempt, thus, $a_{i}.success=True$; otherwise, $a_{i}.success=False$. Each agent $a_{i}$ initially has $a_{i}.success=True$.
  • $\bm{a_{i}.skip}$: If agent $a_{i}$ skips a port, then it stores in this parameter. Initially, $a_{i}.skip=-1$
  • $\bm{a_{i}.divide}$: Initially, all agents are together and each agent has $a_{i}.divide=0$. If and when this group of agents divides into two groups for the first time, each agent sets $a_{i}.divide=1$. There is no change in the values of this variable thereafter.
  • $\bm{a_{i}.grp\_label}$: It stores the ID of the group to which agent $a_{i}$ belongs. Initially $a_{i}.grp\_label=10$. If $a_{i}.grp\_label$ ends with 0, it implies $a_{i}$ is a part of $G_{1}$. Else, if $a_{i}.grp\_label$ ends with 1, it implies $a_{i}$ is a part of $G_{2}$. Later we elaborate on how this variable gets updated.
  • $\bm{a_{i}.count_{1}}$: If agent $a_{i}\in G_{1}$, then it stores the number of consecutive odd rounds with $a_{i}.success=False$.
  • $\bm{a_{i}.count_{2}}$: If agent $a_{i}\in G_{2}$ then it stores the number of consecutive odd rounds with $a_{i}.success=False$.
  • $\bm{a_{i}.count_{3}}$: If agent $a_{i}\in G_{2}$, then it stores the number of new DFS traversals it started from the time $a_{i}$ became a part of this group.
Let $a_{i}$ be a settled agent at node $v$. Any unsettled agent belongs to either $G_{1}$ or $G_{2}$. When an unsettled agent $a_{j}$ while doing its movement (along with other unsettled agents, if any, in its group) according to our algorithm, reaches a node $v$ where an agent $a_{i}$ is already settled, $a_{j}$ needs some information regarding $v$ from $a_{i}$ to continue its traversal. In this regard, each settled agent maintains two parameters, one w.r.t. each of the groups as mentioned below.
[…]
Suppose $n+1$ agents are co-located at a node say $v$ of the graph (rooted configuration). Initially, for each agent $a_{i}$, $a_{i}.grp\_label=10$. Therefore, all agents are a part of $G_{1}$. According to our model, an agent cannot sense the presence of a missing edge unless it attempts to move through it. To tackle this, we take two consecutive rounds to perform one edge traversal. In every odd round, agents move, and in the next even round, it realizes the movement is successful or unsuccessful. Precisely, say an agent $a_{i}$ attempts to move in an odd round $r$. If its movement is successful, it updates its parameter $a_{i}.success$ to $True$ in the even round $r+1$. Otherwise, it updates $a_{i}.success=False$ in the even round $r+1$. Now we proceed with a detailed description of our algorithm.
Algorithm for unsettled agent $\bm{a_{i}}$ of $\bm{G_{1}}$ at odd round: In any odd round, let agent $a_{i}\in G_{1}$ be at node $u$ and $a_{i}.settled=0$. It follows the following steps.
  • $\bm{a_{i}.divide=0}$: Suppose, agent $a_{i}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. Else if agent $a_{i}$ is not the minimum ID, then it moves according to the DFS traversal algorithm. If agent $a_{i}$ reaches a node $v$ and finds an agent with $a_{j}$ with $a_{j}.settled=1$, then it $a_{i}$ does not settle at node $v$ and moves according to the DFS traversal algorithm. In this way, the agent $a_{i}\in G_{1}$ continues the DFS traversal of the graph unless $a_{i}$ encounters a missing edge for the first time. In particular, at some odd round $r$, agent $a_{i}$ moves via an edge $e$, and the movement is not successful at the end of round $r$. This is understood by agent $a_{i}$ in the round $r+1$. In round $r+2$, agent $a_{i}$ updates $divide$ by 1 and divides into two groups using the $Group\_divide$ procedure as follows.
[…]
  • $\bm{a_{i}.divide=1}$: Suppose, agent $a_{i}\in G_{1}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{2}$ at node $w$, then $a_{i}$ updates $a_{i}^{u}(G_{2})$ to $(a_{j}.prt\_in,a_{j}.grp\_label,a_{j}.dfs\_label)$. If $a_{i}\in G_{1}$ reaches a node $w$ and find agent $a_{j}$ with $a_{j}.settled=1$, then it does not settle at the node $v$ and it can understand whether node $w$ is the new node according to the current DFS or previously visited node by comparing $label$, $dfs\_label$ components of $a_{j}^{w}(G_{1})$ with $a_{i}.grp\_label$, $a_{i}.dfs\_label$ respectively. If both matches then node $w$ is visited previously in the current DFS traversal of agent $a_{i}$. Otherwise, if any of those two components do not match, then agent $a_{i}\in G_{1}$ treats $w$ as an unvisited node w.r.t. the current DFS. Round $r+2$ onwards, $a_{i}$ updates its $count_{1}$ parameter as follows. At odd round $r_{1}\geq r+2$, if it finds $a_{i}.success=True$, i.e. it realized a successful movement in the even round $r_{1}-1$, agent $a_{i}$ updates $a_{i}.count_{1}=0$; else if $a_{i}.success=False$, agent $a_{i}$ updates $a_{i}.count_{1}=a_{i}.count_{1}+1$. If at some odd round, $a_{i}$ finds $a_{i}.count_{1}=16n^{2}$, then unsettled agents at node $u$ divide into two groups again using aforementioned $Group\_divide$ procedure. That is, $G_{1}$ divides into two groups only after getting stuck at a node for a particular port for $16n^{2}$ many odd rounds.
Algorithm for unsettled agent $\bm{a_{i}}$ of $\bm{G_{2}}$ at odd round: In any odd round, let agent $a_{i}\in G_{2}$ be at node $u$ and $a_{i}.settled=0$. It follows the following steps.
  • (1)
    $\bm{a_{i}.state=explore}$:
    • (A)
      $\bm{a_{i}.success=True}$: It updates $a_{i}.count_{2}=0$ and does the following.
      • (a)
        Node $\bm{u}$ has no agent with $\bm{setlled=1}$: If $a_{i}$ is the minimum ID at node $u$, it settles at node $u$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ and $a_{i}^{u}(G_{2})$ to $(a_{i}.prt\_in,a_{i}.grp\_label,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{1}$ at node $u$, then $a_{i}$ updates $a_{i}^{u}(G_{1})$ to $(a_{j}.prt\_in,a_{j}.grp\_\\ label,a_{j}.dfs\_label)$. If $a_{i}$ is not the minimum ID, then it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
      • (b)
        Node $\bm{u}$ has an agent with $\bm{setlled=1}$: Let $a_{j}$ be at node $u$ with $a_{j}.settled=1$. If $a_{i}.grp\_label$, $a_{i}.dfs\_label$ matches $label$, $dfs\_label$ components of $a_{j}^{u}(G_{2})$ respectively, then it is a previously visited node. In this case $a_{i}$ set $a_{i}.state=backtrack$, $a_{i}.prt\_out=a_{i}.prt\_in$, and it moves through $a_{i}.prt\_out$. Otherwise, it is an unvisited node w.r.t. the current DFS of $a_{i}$. In this case, it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
    • (B)
      $\bm{a_{i}.success=False}$: It does the following.
      • (a)
        There is an agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: Let $a_{j}$ be settled at node $u$, $b_{i}\in G_{1}$ be at node $u$, and $b_{i}.settled=0$. In this case, there are two cases.
        • (i)
          If $a_{i}.prt\_out=b_{i}.prt\_out$ (observe that based on communication, agent $a_{i}$ can compute what agent $b_{i}$ computes, and vice versa), then there are two cases possible.
          ($\alpha$) $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is $\bm{-1}$: If $a_{i}.prt\_out=\delta_{u}-1$, then agent $a_{i}$ sets $a_{i}.state=explore$, stores $a_{i}.skip=-1$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=0$, and moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out\neq\delta_{u}-1$, then it sets $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{u}$ and sets $a_{i}.count_{2}=0$ and move through $a_{i}.prt\_out$.
          ($\beta$) $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is not $\bm{-1}$: It sets $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{u}$ and sets $a_{i}.count_{2}=0$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
        • (ii)
          If $a_{i}.prt\_out\neq b_{i}.prt\_out$, it updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. In this case, agent $a_{i}$ sets $a_{i}.state=explore$, stores $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Note that $a_{i}.count_{3}$ always remains less than $4n^{2}$ in this case as justified in the correctness.
      • (b)
        There is no agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: It updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Else, if $a_{i}.count_{3}=4n^{2}$, it divides into two groups using $Group\_divide$ procedure mentioned in the algorithm for unsettled agents of $G_{1}$. Note that if $a_{i}.count_{3}=4n^{2}$, then $G_{1}$ is already dispersed as justified in the correctness.
  • (2)
    $\bm{a_{i}.state=backtrack}$:
    • (A)
      If $\bm{a_{i}.success=True}:$ It updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$ and $a_{i}.count_{2}=0$. Let $a_{j}$ be the agent at node $u$ with $a_{j}.settled=1$. Based on $parent$ component of $a_{j}^{u}(G_{2})$, there are two cases.
      • (a)
        $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is $\bm{-1}$: If agent $a_{i}$ finds $parent$ component of $a_{j}^{u}(G_{2})$ to be $-1$, then node $u$ is the root node of the current DFS traversal of agent $a_{i}$. In this case, if $a_{i}.prt\_out=$ minimum available port except for $a_{i}.skip$, then it sets $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, and moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out\neq$ minimum available port except for $a_{i}.skip$, then it does the following.
        If $a_{i}.prt\_out=a_{i}.skip$, then $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{v}$. If $a_{i}.prt\_out=$ minimum available port except for $a_{i}.skip$, then it sets $a_{i}.state=explore$, $a_{i}.skip=-1$, $a_{i}.prt\_out=0$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, and moves through $a_{i}.prt\_out$. Else if, it sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
        If $a_{i}.prt\_out\neq a_{i}.skip$, it sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
      • (b)
        $\bm{parent}$ component of $\bm{a_{j}^{u}(G_{2})}$ is not $\bm{-1}$: If agent $a_{i}$ does not find $parent$ component of $a_{j}^{u}(G_{2})$ to be $-1$, then node $u$ is not the root node of the current DFS traversal of agent $a_{i}$. If $a_{i}.prt\_out$ matches with $parent$ component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ moves through $a_{i}.prt\_out$. Else if $a_{i}.prt\_out$ does not match with $parent$ component of $a_{j}^{u}(G_{2})$, then agent $a_{i}$ sets $a_{i}.state=explore$ and moves through $a_{i}.prt\_out$.
    • (B)
      If $\bm{a_{i}.success=False}:$ Two cases are possible.
      • (a)
        There is an agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: Let $a_{j}$ be settled at node $u$, $b_{i}\in G_{1}$ be at node $u$, and $b_{i}.settled=0$. In this case, there are two cases.
        • (i)
          If $a_{i}.prt\_out=b_{i}.prt\_out$ (observe that based on communication, agent $a_{i}$ can compute what agent $b_{i}$ computes, and vice versa), then agent $a_{i}$ does the following. It updates $a_{i}.count_{2}=0$ and sets $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$,
          $a_{i}.dfs\_label=a_{i}.dfs\_label+1$. Agent $a_{i}$ moves through $a_{i}.prt\_out$.
        • (ii)
          If $a_{i}.prt\_out\neq b_{i}.prt\_out$, it updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.port=a_{i}.prt\_out$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.prt\_out=0$, $a_{i}.skip=-1$, and moves through $a_{i}.prt\_out$. Note that $a_{i}.count_{3}$ always remains less than $4n^{2}$ in this case as justified in the correctness.
      • (b)
        There is no agent from $\bm{G_{1}}$ with $\bm{settled=0}$ at node $\bm{u}$: It updates $a_{i}.count_{2}=a_{i}.count_{2}+1$. If $a_{i}.count_{2}<4n^{2}$, then move through $a_{i}.prt\_out$. Else, if $a_{i}.count_{2}=4n^{2}$, then $a_{i}$ updates $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$. If $a_{i}.count_{3}<4n^{2}$, then agent $a_{i}$ sets $a_{i}.state=explore$, $a_{i}.dfs\_label=a_{i}.dfs\_label+1$, $a_{i}.skip=a_{i}.prt\_out$, $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$, and moves through $a_{i}.prt\_out$. Else, if $a_{i}.count_{3}=4n^{2}$, it divides into two groups using $Group\_divide$ procedure mentioned in the algorithm for unsettled agents of $G_{1}$.
Algorithm for unsettled agent $\bm{a_{i}}$ of at even round: At even round, each unsettled agent $a_{i}$ of $G_{1}$ or $G_{2}$ updates $a_{i}.success$ parameter. If their movement in the previous rounds is successful, agent $a_{i}$ updates their $a_{i}.success=True$. Otherwise, it updates their parameter $a_{i}.success=False$.
Algorithm for settled agents: Let $a_{i}$ be at node $u$ with $a_{i}.settled=1$. It has two parameter $a_{i}^{u}(G_{1})$ and $a_{i}^{u}(G_{2})$. At each odd round based on the communication, it computes the decision of $G_{1}$ and $G_{2}$, and it stores $G_{1}$ information in $a_{i}^{u}(G_{1})$ and $G_{2}$ information in $a_{i}^{u}(G_{2})$. In move, it does not do anything. At even round, it does not do anything.
The pseudocode of our algorithm is provided in Section 6.
3.2 Correctness and analysis of algorithm
Let $m$ be the number of edges in $G$. At the beginning of the algorithm, $n+1$ agents are co-located at a node of $G$. Therefore, all agents can learn the value of $n$. Initially, each agent $a_{i}$ is part of $G_{1}$ as $a_{i}.grp\_label=10$. Assume that the adversary does not remove an edge in the first $4m$ rounds. In this case, all agents reach the dispersion configuration because this is nothing but the dispersion algorithm using the DFS traversal of [3]. Hence, all agents reach dispersion configuration in the first $4m$ rounds.
Let in round $r<4m$, all agents in $G_{1}$ be at node $u$ and want to go through edge $e=(u,\,v)$. At the round $r$, if the adversary removes an edge $e$, then the movement of agents in $G_{1}$ becomes unsuccessful. In this case, we divide the group $G_{1}$ into two new groups $G_{1}$ and $G_{2}$. And, both the groups consider the current node $u$ as the root node for the new DFS traversal and restart the algorithm for DFS via port 0. Let $r^{\prime}\geq r$ be round when agents in $G_{1}$ and $G_{2}$ are at node $u^{\prime}$ and want to go through edge $e^{\prime}=(u^{\prime},\,v^{\prime})$ via port $p$, and their movement becomes unsuccessful due to the deletion of the edge $e^{\prime}$ by the adversary. We have the following claim.
Claim 1: If edge $e^{\prime}$ does not appear within next $16n^{2}$ odd rounds after it is deleted by the adversary in round $r^{\prime}$, then agents in $G_{2}$ get dispersed in those $16n^{2}$ rounds.
Proof
In round $r^{\prime}$, for $a_{i}\in G_{2}$, $a_{i}.state$ is either $explore$ or $backtrack$. Note that each unsettled agent in $G_{2}$ has the same state in any round. Each agent $a_{i}$ takes $4m$ odd rounds to reach node $v^{\prime}$ and tries to move through edge $(v^{\prime},\,u^{\prime})$ as justified below.
  • if $\bm{a_{i}.state=backtrack}$: If agent $a_{i}\in G_{2}$ is with $a_{i}.state=backtrack$, then it starts a new DFS considering $u^{\prime}$ as a new root node and stores $a_{i}.skip=p$. In the next $4m$ odd rounds, it runs the DFS traversal on $G-\{e^{\prime}\}$. Therefore, it reaches node $v^{\prime}$ unless all agents of $G_{2}$ get dispersed.
  • if $\bm{a_{i}.state=explore}$: In this case, it continues its current DFS traversal by setting $a_{i}.prt\_out=(p+1)$ mod $\delta_{u^{\prime}}$. There are two cases: Case (i) the node $v^{\prime}$ is already visited node in its current DFS traversal by the round $r^{\prime}$, or Case (ii) the node $v^{\prime}$ is not visited in its current DFS traversal by the round $r^{\prime}$.
    Case (i): Let $a_{i}$ has taken $t$ odd rounds to visit node $v^{\prime}$ in the current DFS before trying to traverse through edge $e^{\prime}$. After round $r^{\prime}$, agent $a_{i}$ continues the DFS traversal via $(p+1)$ mod $\delta_{u^{\prime}}$ port. It may be possible that it does not reach node $v^{\prime}$ via the remaining part of the DFS traversal because it was unable to move by port $p$. According to our algorithm, agent $a_{i}\in G_{2}$ reaches the root node (say $w$) of the current DFS within $4m-t$ odd rounds (as to complete the current agents need $4m$ odd rounds) and, if there is some unsettled agent $a_{i}\in G_{2}$ at node $w$, it starts a new DFS by increasing $a_{i}.dfs\_label$ by 1, considering $w$ as root node and takes at most $t$ rounds to reach node $v^{\prime}$. In total, agent $a_{i}\in G_{2}$ takes $4m-t+t$=$4m$ odd rounds to reach node $v^{\prime}$.
    Case (ii): Suppose, the current DFS traversal of agent $a_{i}$ has been executed for $t^{\prime}$ odd rounds. To complete the current DFS, agent $a_{i}\in G_{2}$ takes $4m-t^{\prime}$ odd rounds more. In this case, agent $a_{i}$ reaches node $v^{\prime}$ within $4m-t^{\prime}$ odd rounds.
If at least one unsettled agent in $G_{2}$ is at node $v^{\prime}$, each agent $a_{i}\in G_{2}$ tries to move edge $(v^{\prime},\,u^{\prime})$. Whenever the movement is unsuccessful each agent $a_{i}\in G_{2}$ increases $a_{i}.count_{2}$ by 1. Since $e^{\prime}$ is missing for $16n^{2}$ consecutive odd rounds from round $r^{\prime}$, therefore for each agent $a_{i}\in G_{2}$, $a_{i}.count_{2}$ becomes $4n^{2}$ as from round $r^{\prime}$ to reach node $v^{\prime}$ it takes $4m$ odd round. In total time from $r^{\prime}$ passes by is $4m+4n^{2}<16n^{2}$ odd rounds. Therefore, $a_{i}.count_{2}$ reaches $4n^{2}$. In this case, agent $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$\&& $a_{i}.dfs\_label$ by 1, considers node $v^{\prime}$ the root and starts the new DFS from node $v^{\prime}$ as well as stores the outgoing port of edge $(v^{\prime},\,u^{\prime})$ as $a_{i}.skip$. In the next $4m\leq 4n^{2}$ odd rounds, each agent $a_{i}\in G_{2}$ completes the DFS traversal on $G-e^{\prime}$. It is because even if it reaches node $u^{\prime}$ and wants to go through edge $e^{\prime}$ via port $p$, it reaches as $a_{i}.state=explore$. According to our algorithm, agent $a_{i}$ changes the outgoing port by $(p+1)$ mode $\delta_{u}$ and continues its DFS traversal on $G-e^{\prime}$. The total odd rounds till dispersion of agents in $G_{2}$ is at most $12n^{2}$ odd rounds. Therefore, if edge $e^{\prime}$ does not appear in consecutive $16n^{2}$ rounds (we explain later why we are using $16n^{2}$ instead of $12n^{2}$), then the group $G_{2}$ is successfully dispersed in $G$.∎
If the edge $e^{\prime}$ reappears at some round before the wait time of the group $G_{1}$, then agents in $G_{1}$ move at least one round of its current DFS. Now we make a claim that guarantees the dispersion of agents in $G_{1}$.
Claim 2: For any unsettled agent $a_{i}\in G_{2}$, if $a_{i}.count_{3}$ matches with $4n^{2}$, then agents in $G_{1}$ are already dispersed.
Proof
Let $T$ be an odd round when edge $e^{\prime}$ reappears and agents in $G_{1}$ move at least one round on its current DFS traversal. Note that $T>r^{\prime}$. After round $T$ round onwards the following cases are possible.
  • Case (A) Agents in group $G_{1}$ and $G_{2}$ want to go through the same deleted edge: In this case, the group $G_{2}$ meets with $G_{1}$ at node $u_{1}$ and both want to go through the same edge $e_{1}=(u_{1},\,v_{1})$ in round $r_{1}$ (say corresponding port $p_{1}$ of $u_{1}$). There are two scenarios: (i) $G_{1}$ reaches node $u_{1}$ in round $r^{\prime\prime}<r_{1}$ but $G_{2}$ is not present, or (ii) $G_{2}$ reaches node $u_{1}$ in round $r^{\prime\prime\prime}<r_{1}$ but $G_{1}$ is not present in round $r^{\prime\prime\prime}$. In both cases, edge $e_{1}$ is deleted from round $r^{\prime\prime}$ or $r^{\prime\prime\prime}$ onwards (note that $r^{\prime\prime},r^{\prime\prime\prime}\geq T$). Within the next $4m$ odd rounds, the other group $G_{2}$ may come to node $u_{1}$ and move through edge $e_{1}$ (this is similar to Case (i) &\&& (ii) as discussed in Claim 1). These two situations change to the case that we have discussed in Claim 1. After round $r_{1}$, if the edge is deleted for $12n^{2}$ odd rounds consecutively, then $G_{2}$ gets dispersed. Due to this reason, we are giving waiting time $16n^{2}$ instead of $12n^{2}$ to $G_{1}$.
  • Case (B) Agents in $G_{2}$ want to go through deleted edge: In this case, agents in $G_{2}$ at node $u_{2}$ try to move via edge $e_{2}=(u_{2},\,v_{2})$ using port $p_{2}$. Agent $a_{i}\in G_{2}$ has parameter $a_{i}.count_{2}$ and it is increasing this parameter by 1 whenever its movement becomes unsuccessful. If its movement is successful, then $a_{i}.count_{2}=0$. If at some odd round, say $t^{\prime}$, $a_{i}.count_{2}$ reaches to $4n^{2}$ (it means $4n^{2}$ consecutive odd rounds edge $e^{\prime}$ is missing), then agent $a_{i}\in G_{2}$ restarts the new DFS and stores $a_{i}.skip=p_{2}$ corresponding edge $e^{\prime}$ through which it is trying to move. It is important to note that if edge $e_{2}$ is missing for consecutive $4n^{2}$ odd rounds, then two cases are possible which are as follows.
    • Case (a): While $G_{2}$ is waiting for $e_{2}$ to appear, then within $4n^{2}$ odd rounds unsettled agents in $G_{1}$ get stuck at node $u_{2}$ (or $v_{2}$). If each unsettled agent $a_{j}\in G_{1}$ reaches node $u_{2}$ and wants to go through edge $e_{2}$, then it is similar to Case (A). If each unsettled agent $a_{j}\in G_{1}$ gets stuck at node $v_{2}$, then it tries to move through edge $(v_{2},\,u_{2})$ and starts increasing $a_{j}.count_{1}$ parameter by 1 in each unsuccessful movement. When $a_{i}.count_{2}=4n^{2}$, then agent $a_{i}\in G_{2}$ restarts new DFS, stores $a_{i}.skip=p_{2}$ and increases parameter $a_{i}.count_{3}$ by one. If edge $e_{2}$ does not appear next odd $4m$ rounds (note that the parameter for agent $a_{j}\in G_{1}$ is less than $16n^{2}$), then agents in $G_{2}$ achieve the dispersion. It is because even if it reaches node $v_{2}$ and wants to go through edge $(v_{2},\,u_{2})$ via some port $p$, it reaches as $a_{i}.state=explore$. Agent $a_{i}$ changes the outgoing port by $(p+1)$ mode $\delta_{v_{2}}$ and continues its DFS traversal on $G-\{e_{2}\}$. Therefore, if edge $e_{2}$ does not appear in the next consecutive $4m$ odd rounds, then the group $G_{2}$ is successfully dispersed at the nodes of $G$. If edge $e_{2}$ reappears then agents in $G_{1}$ move at least one round of its current DFS.
    • Case (b): The DFS traversal of $G_{1}$ is unaffected due to the missing edge $e_{2}$. In this case, agents in $G_{1}$ get dispersed.
It is important to note that if unsettled agent $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$, then agent $a_{j}\in G_{1}$ moves at least one round of its current DFS. Therefore, whenever $a_{i}.count_{3}$ reaches $4n^{2}$, agents in $G_{2}$ can understand that agents in $G_{1}$ are dispersed as $G_{1}$ already ran its dispersion algorithm via DFS for at least $4m$ odd rounds without deviating from the path. ∎
Based on Claim 1 &\&& 2, we have the following lemma.
Lemma 2
Our algorithm correctly solves $(n+1)$-Dispersion on TVG for co-located $n+1$ agents.
Proof
Due to Claim 1 &\&& 2, we can say either $G_{1}$ or $G_{2}$ is dispersed and unsettled agents in $G_{1}\,(G_{2})$ understand whether agents in $G_{2}\,(G_{1})$ are dispersed. And, agents in $G_{1}\,($or $G_{2})$ divide into two new groups. In this way, in the end, group $G_{1}$ (and $G_{2}$) contains 1 agent. Both groups start running our algorithm, and due to Claim 1 &\&& 2, we can say that either agent in $G_{1}$ or $G_{2}$ is settled at some node. Therefore, our algorithm correctly solves $(n+1)$-Dispersion on TVG for co-located $n+1$ agents. ∎
Due to Lemma 2, we can say that one group is divided into two groups only after the other group achieves dispersion. Since we have $n+1$ agents, in this way, all agents are getting dispersed except one agent which will be part of the last remaining group.
Lemma 3
The time complexity of our algorithm is $O(n^{4}\cdot\log n)$.
Proof
In Lemma 2, we have shown that the agents achieve the dispersion. If the adversary does not remove any edge in the first $4m$ odd rounds then all agents are dispersed except one agent. In odd round $r$ (where $r<4m$), if the adversary removes the edge $e$ and the group of agents wants to go through the edge $e$ according to their DFS algorithm, then they divide into two groups. We show that in the next $32n^{4}$ odd rounds, either $G_{1}$ or $G_{2}$ gets dispersed. We have proved that whenever $a_{i}\in G_{2}$ increases $a_{i}.count_{3}$ by 1, then agents in $G_{1}$ move at least one round of its DFS or agents in $G_{1}$ get dispersed. Therefore, if agents in $G_{2}$ are not dispersed and for $a_{i}\in G_{2}$, $a_{i}.count_{3}$ reaches $4n^{2}$, then agents in $G_{2}$ can understand that agents in $G_{1}$ are dispersed. Suppose agents $G_{2}$ runs each DFS algorithm for $4m$ rounds, therefore $4(m+n^{2})\cdot 4n^{2}\leq 32n^{4}$ odd rounds are required to disperse $G_{1}$. We also have shown if $G_{1}$ waits for an edge to reappear for $16n^{2}$ rounds, then $G_{2}$ is dispersed. Therefore, after dividing into two groups, in the next $32n^{4}$ odd rounds, either $G_{1}$ or $G_{2}$ is dispersed. Therefore, either $G_{1}$ or $G_{2}$ gets dispersed in the first $32n^{4}+4m$ odd rounds. Including even rounds, we can say that after dividing into two groups, in the next $64n^{4}+8m$ rounds, either $G_{1}$ or $G_{2}$ is dispersed.
Using Lemma 2, if $G_{1}$ (or $G_{2}$) understands that $G_{2}$ (or $G_{1}$) is dispersed, then $G_{1}$ (or $G_{2}$) divides into two new groups, and both groups contain half agents of $G_{1}$ (or $G_{2}$). These two groups repeat the same procedure. The number of such group divisions is $O(\log n)$, and one of two groups is dispersed in $64n^{4}+8m$ rounds. Therefore, our algorithm takes $O(n^{4}\cdot\log n)$ rounds to solve $(n+1)$-Dispersion on TVG for co-located $n+1$ agents. ∎
Lemma 4
In our algorithm, agents require $\Theta(\log n)$ memory.
Proof
For agent $a_{i}$, $a_{i}.grp\_label$ is nothing but a binary string. Whenever we divide the group, the $a_{i}.grp\_label$ is appended by 0 or 1. The number of group divisions is $O(\log n)$. Therefore, the length of $a_{i}.grp\_label$ is $O(\log n)$, which agent $a_{i}$ can store in its $O(\log n)$ memory. According to Lemma 3, all agents achieve dispersion in $O(n^{4}\cdot\log n)$ rounds. Therefore, $a_{i}.dfs\_label$ can not be more than $O(n^{4}\cdot\log n)$ which agent $a_{i}$ can store in its $O(\log n)$ memory. Apart from these two parameters, all unsettled agents $a_{i}$ either store boolean variables or port information. Such parameters are constant as per our algorithm. To store each parameter, we need $O(\log n)$ storage. Therefore, each unsettled agent $a_{i}$ needs $O(\log n)$ memory. All settled agents maintain two parameters as per our algorithm $a_{i}^{u}(G_{1})$ and $a_{i}^{u}(G_{2})$. To store each parameter, settled agents need $O(\log n)$ memory. Therefore, each settled agent $a_{i}$ needs $O(\log n)$ memory.
Due to Lemma 1, the memory lower bound for the dispersion of $k\geq n+1$ agents is $\Omega(\log n)$, and in our algorithm agents use $O(\log n)$ memory each. Therefore, the agents in our algorithm require $\Theta(\log n)$ memory due to $k=n+1$. ∎
Now, we are ready to provide our main theorem.
Theorem 3.1
Our algorithm solves $k$-Dispersion on TVG for $k=n+1$ co-located agents in $\mathcal{G}$ in $O(n^{4}\cdot\log n)$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The proof follows from Lemma 2, 3, and 4. ∎
Observation 1
All agents know the value of $n$ and $m\leq n^{2}$. If all agents want to achieve the termination, can calculate the value of $72n^{4}\cdot\lceil\log n\rceil$ and terminate after $72n^{4}\cdot\lceil\log n\rceil$ rounds.
4 Discussion
In this section, we give a discussion of how to solve $(pn+1)$-Dispersion on TVG without global communication using 0-hop visibility. We also provide an algorithm to solve $(pn)$-Dispersion on TVG with global communication and 1-hop visibility using the algorithm of [16].
4.1 $(pn+1)$-Dispersion on TVG
In this section, we demonstrate how to extend the algorithm from Section 3 to address the $(pn+1)$-Dispersion on TVGs, where $p>1$. In the dispersed configuration, each node contains $p$ agents, except one node contains $p+1$ agents. We assume that agents are aware of the parameter $n$, and $pn+1$ agents are co-located at some node $v$. In the first round, agents distribute into $n$ groups, say $g_{1}$, $g_{2}$, …, $g_{n}$. Since agents are co-located, they can compute the value of $p$ using $n$. The number of agents in each $g_{i}$, $11\leq i\leq n-$ 1, is $p$, and $g_{n}$ is $p+1$. Each agent $a_{i}$ in $g_{j}$ remembers $j$ in $a_{i}.ID_{1}$ parameter, where $1\leq j\leq n$. In the same manner, agents start executing the algorithm mentioned in Section 3. In this algorithm, whenever agents find a node without any agents, then $g_{i}$ gets settled in place of an agent. Here, the remaining unsettled $g_{i}$s divide into two groups, say $G_{1}$ and $G_{2}$, in place of dividing groups of agents into 2 groups like Section 3. In Lemma 2, we have shown one of $G_{1}$ or $G_{2}$ gets dispersed based on their parameters. When $G_{1}$ ($G_{2}$) is left with one $g_{j}$ (unsettled group), and $G_{1}$ ($G_{2}$) is sure that $G_{2}$ ($G_{1}$) has been dispersed, then they follow the following steps. W.l.o.g., let $G_{1}$ be with one $g_{i}$ at node $v$, and $G_{2}$ has been dispersed. In this case, it is important to note that 1 node is left, which has no agents. If $v$ is such a node, then dispersion has been achieved. If $v$ is not such a node, then there is some node $u$ which has no agent. In our algorithm mentioned in Section 3, there were two agents left at the end who were searching for the single empty node via different paths, one gets to that node as the adversary can not stop both at a round, and the other settled when it is sure that it the only one that is left to settle. Here, we can not use our algorithm idea because the adversary can always stop one group, i.e., the last remaining group, to find the single empty node in the graph. Below we provide the idea with correctness how the group $g_{i}$ gets settled at node $u$. There are two cases, and we discuss each case as follows.
  1. 1.
    Case $\bm{|g_{i}|=p+1}$ : The agents start running the algorithm mentioned in Section 3 in the following manner. Now $g_{i}$ executes DFS for the single node, say $u$, that has less than $p$ agents, actually no agents at this point of time, on it. Whenever the missing edge appears, $g_{i}$ divides into two groups, say $G_{1}$ and $G_{2}$, and starts executing DFS. As per Lemma 2, one of the groups $G_{1}$ or $G_{2}$ finds the node $u$ that has $<p$ agents and settles there. This continues until $u$ has at least $p$ settled agents. In the end, like our algorithm, when $G_{1}$ ($G_{2}$) is left with one agent, and it is sure that $G_{2}$ ($G_{1}$) has been dispersed, it settles at the current node, say $v$, as by this time the remaining $p$ agents of $g_{i}$ must have settled at $u$. Note that $v$ can be the node $u$ as well, but irrespective of what $v$ is, the dispersed configuration is achieved as exactly one node ($v$) contains $p+1$ agents.
  2. 2.
    Case $\bm{|g_{i}|=p}$ : In this case, $g_{j}$ with $p+1$ has been settled at some node $u_{1}$ in the earlier round. In observation 1, we have given the number of rounds of termination. In this case, agents in $g_{i}$ wait till that round. Agents in $g_{i}$ become group $G_{1}$, and the largest ID at $g_{j}$ becomes $G_{2}$. They start executing DFS as per our algorithm. As per Claim 1, if $G_{1}$ waits for consecutive $16n^{2}$ odd rounds, then $G_{2}$ visits each node at least once. In this case, $G_{2}$ and $G_{1}$ meet, and they become a group of $p+1$ agents, which is an earlier case. If $G_{1}$ does not wait for consecutive $16n^{2}$ odd rounds, then it is moving on the current DFS. And in at most $16n^{2}\times 4n^{2}$ odd rounds, it reaches node $u$ and settles. In at most $64n^{2}$ odd rounds, the agent in $G_{2}$ becomes settled at its current node. This ensures that the agents achieve a dispersed configuration because, after $64n^{2}$ odd rounds, if $G_{2}$ has not met $G_{1}$, it indicates that $G_{1}$ has completed its depth-first search (DFS).
Time Complexity of solving $(pn+1)$-Dispersion on TVG:
Now we are ready to state the final theorem.
Theorem 4.1
$(pn+1)$-Dispersion on TVG can be solved in $O(n^{4}\cdot max(\log n,\,\log p))$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The correctness is justified in the above discussion. Since agents are using the algorithm mentioned in Section 3. Therefore, $\Theta(\log n)$ memory per agent is required. Since $n$ many $g_{i}$s are executing the algorithm, therefore, in $O(n^{4}\cdot\log n)$ rounds, one of the aforementioned cases is possible. In the first case, agents take $O(n^{4}\cdot\log p)$ extra rounds to achieve the dispersed configuration, and in the second case if both groups meet with each other within $64n^{2}$ odd rounds, then they convert into the first case, and take $O(n^{4}\cdot\log p)$ extra rounds to achieve the dispersed configuration. In the second case, if they do not meet with each other, then $g_{i}$ has reached node $u$ within $64n^{2}$, and the dispersion has been achieved. Therefore, the total time is $O(n^{4}\cdot\log n)+O(n^{4}\cdot\log p)=O(n^{4}\cdot max(\log n,\,\log p))$. This completes the proof. ∎
4.2 $k$-Dispersion on TVG
In Section 2, we have shown that to solve $(pn)$-Dispersion on TVG, agents need 1-hop visibility and global communication. In this section, we will extend the algorithm of [16] to solve $k$-Dispersion on TVG when agents are equipped with 1-hop visibility and global communication. Recall that their algorithm works in the weak dynamic graph model in comparison to TVG, i.e., the adversary can delete or add edges arbitrarily, maintaining the graph in each round connected. Let’s refer to the algorithm as $Algo\_weak\_Disp(k)$. First, we recall a few results from their work.
Theorem 4.2
[16] Given $k\leq n$ agents placed arbitrarily on the nodes of any $n$-node graph $\mathcal{G}_{r}$, where $\mathcal{G}_{r}$ is the graph at round $r\geq 0$. $Algo\_weak\_Disp(k)$ solves the dispersion in $\Theta(k)$ rounds with $\Theta(\log k)$ bits at each agent in the synchronous setting with global communication and 1-hop visibility. Also, all agents understand that dispersion has been achieved and terminated.
Based on $Algo\_weak\_Disp(k)$, the authors have the following lemma.
Lemma 5
[16] Consider any $n$-node dynamic graph $\mathcal{G}_{r}$ at round $r\geq 0$. If $k\leq n$ agents are positioned on $l$ nodes of $\mathcal{G}_{r}$ at the beginning of round $r$, then at the beginning of round $r+1$, the agents are positioned on at least $l+1$ nodes of $\mathcal{G}_{r}$.
Now, we are ready to provide how we can extend $Algo\_weak\_Disp(k)$ to solve $k$-Dispersion on TVG when $k$ agents in $G$ need 1-hop visibility and global communication. Note that this idea works even if agents are scattered, agents do not know $n$, and the adversary deletes more than 1 edge while maintaining $\mathcal{G}$ connected.
The agents run $Algo\_weak\_Disp(k)$. Using Lemma 5, we can observe that the number of nodes without any agents is reduced by at least 1 in each round. Therefore, after $n$ rounds, all nodes have at least 1 agent. At each round $r$, agents check whether their neighbour is occupied by some agent or not and share this information using global communication with all agents. At some round $r$, when agents do not get such information, agents can understand that each node has at least 1 agent. At every round $r^{\prime}>r$, at each node $v$, the largest ID agent becomes a leader, and it shares the information of the number of agents at node $v$ along with its ID. In this way, at every round $r^{\prime}$ with the help of global communication, agents can learn the value of $k$ using the number of agents at other nodes, and they can also learn the value of $n$ based on the number of distinct IDs they receive. Therefore, they do not have to remember $k$ and $n$, and in each round $r^{\prime}$, they can learn $k$ and $n$. This information allows agents to compute $\lceil k/n\rceil$. From round $r+2$, they start executing $Algo\_weak\_Disp(k)$. In each round, agents can learn $\lceil k/n\rceil$. This is important otherwise the memory requirement per agent will be $O(max(\log p,\log n))$. When they find at least one node has less than $\lceil k/n\rceil$ agents. They can follow $Algo\_weak\_Disp(k)$. Using Lemma 5, we can say that if there is a node $u$ with more than $\lceil k/n\rceil$ agents, then the extra agent from node $u$ moves to the other node $v$ which has less than $\lceil k/n\rceil$ agents. In this way, they can do this in each round. When agents find no node with more than $\lceil k/n\rceil$ agents, they can understand $k$-Dispersion on TVG has been achieved. Since $k$ agents are present, agents can achieve $k$-Dispersion on TVG in $O(k)$ rounds. In each round, agents need to remember their ID, and it does not have to remember $k$ or $n$. Therefore, agents need $O(\log n)$ memory as ID range is $[1,\,n^{c}]$, where $c$ is some constant. Using Lemma 1, we can say $\Theta(\log n)$ memory per agent is required. Based on this, we have the final theorem.
Theorem 4.3
Let $k$ agents be placed arbitrarily on the nodes of $G$, and the adversary can delete edges arbitrarily from footprint $G$, maintaining $\mathcal{G}$ connected. $Algo\_weak\_Disp(k)$ solves $k$-Dispersion on TVG in $O(k)$ rounds with $\Theta(\log n)$ bits of memory at each agent with global communication and 1-hop visibility. Also, all agents understand that dispersion has been achieved and terminated.
5 Conclusion
In this work, we have studied the Dispersion on TVG on $\mathcal{G}$. We found that with $k=n$ agents starting from a rooted configuration, 1-hop visibility and global communication are necessary to solve the Dispersion on TVG. Additionally, we have shown that if $k\geq n+1$, agents do not require 1-hop visibility and global communication to solve the Dispersion on TVG from any rooted initial configuration. One can try to improve the time complexity for the case $k\geq n+1$. It would also be interesting to explore the Dispersion on TVG for $k\geq n+1$ when agents start from arbitrary initial configuration. Another interesting question that remains open is whether global communication and 1-hop visibility are necessary to solve dispersion for $k<n$ agents.
6 Pseudocodes
1 while True do
2       $a_{i}.r=a_{i}.r+1$
3       if $a_{i}$ belongs to $G_{1}$ and $a_{i}.settled=0$ then
4             call Algorithm 3
5      else if $a_{i}$ belongs to $G_{2}$ and $a_{i}.settled=0$ then
6             call Algorithm 2
7      else if $a_{i}.settled=1$, there is $a_{j}$ from $G_{1}$ at the same node and $a_{j}.settled=0$ then
8             call Algorithm 7
9      else if $a_{i}.settled=1$, there is $a_{j}$ from $G_{2}$ at the same node and $a_{j}.settled=0$ then
10             call Algorithm 8
Algorithm 1 Dispersion in $1$-TVG with $n+1$ agents
1 if $a_{.}state=explore$ then
2       call Algorithm 4
3else
4       call Algorithm 5
Algorithm 2 Movement of agent $a_{i}\in G_{2}$ at $v$ with $a_{i}.settled=0$
1 if $a_{i}.state=explore$ then
2       if $r$ mod 2 =1 then
3             if $a_{i}.success=True$ then
4                   set $a_{i}.count_{1}=0$
5                   if there is no agent at node $v$ with $settled=1$ then
6                         if $a_{i}.ID$ is the minimum from all the unsettled agents present at the current node then
7                               set $a_{i}.settled=1$
8                               update $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$
9                               if there is an agent $a_{j}$ from $G_{2}$ with $a_{j}.settled=0$ then
10                                     update $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
11                        else
12                               set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
13                               if $a_{i}.prt\_out=a_{i}.prt\_in$ then
14                                     set $a_{i}.state=backtrack$
15                                     move through $a_{i}.prt\_out$
16                              else
17                                     move through $a_{i}.prt\_out$
18                  else if there is an agent $r_{j}$ at node $v$ with $settled=1$ then
19                         $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{v}$
20                         if $label=a_{i}.grp\_label$, where $label$ is the component of $r_{j}^{v}(G_{1})$ then
21                               $a_{i}.state=backtrack$
22                               move through $a_{i}.prt\_in$
23                        else if $label\neq a_{i}.grp\_label$, where $label$ is the component of $r_{j}^{v}(G_{1})$ then
24                               move through $a_{i}.prt\_out$
25            else if $a_{i}.success=False$ then
26                   if $a_{i}.divide=0$ then
27                         set $a_{i}.divide=1$
28                         call $Group\_divide()$ (Algorithm 6)
29
30                  else if $a_{i}.divide=1$ then
31                         update $a_{i}.count_{1}=a_{i}.count_{1}+1$
32                         if $a_{i}.count_{1}=16n^{2}$ then
33                               call $Group\_divide()$ (Algorithm 6)
34
35                        else if $a_{i}.count_{1}<16n^{2}$ then
36                               move through $a_{i}.prt\_out$
37      else if $r$ mod 2=0 then
38             if the movement of agent $a_{i}$ in round $r-1$ is successful then
39                   set $a_{i}.success=True$
40            else
41                   set $a_{i}.success=False$
42else if $a_{i}.state=backtrack$ then
43       if $r$ mod 2=1 then
44             if $a_{i}.success=True$ then
45                   set $a_{i}.count_{1}=0$
46                   if $parent$ component of $r_{j}^{v}(G_{1})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
47                         set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
48                         set $a_{i}.state=explore$
49                         move through $a_{i}.prt\_out$
50                  else if $parent$ component of $r_{j}^{v}(G_{1})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
51
52                        set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
53                         if $a_{i}.prt\_out=parent$ and $parent$ is the component of $r_{j}^{v}(G_{1})$, where $r_{j}$ is the settled agent at node $v$ then
54                               move through $a_{i}.prt\_out$
55                        else
56                               set $a_{i}.state=explore$
57                               move through $a_{i}.prt\_out$
58            else if $a_{i}.success=False$ then
59                   if $a_{i}.divide=0$ then
60                         set $a_{i}.divide=1$
61                         call $Group\_divide()$ (Algorithm 6)
62
63                  else if $a_{i}.divide=1$ then
64                         update $a_{i}.count_{1}=a_{i}.count_{1}+1$
65                         if $a_{i}.count_{1}=16n^{2}$ then
66                               call $Group\_divide()$ (Algorithm 6)
67
68                        else if $a_{i}.count_{1}<16n^{2}$ then
69                               move through $a_{i}.prt\_out$
70      if $r$ mod 2=0 then
71             if the movement of agent $a_{i}$ in round $r-1$ is successful then
72                   set $a_{i}.success=True$
73            else
74                   set $a_{i}.success=False$
Algorithm 3 Movement of agent $a_{i}\in G_{1}$ at $v$ with $a_{i}.settled=0$
1 if $r$ mod 2 =1 then
2       if $a_{i}.success=True$ then
3             set $a_{i}.count_{2}=0$
4             if there is no agent at node $v$ with $settled=1$ then
5                   if $a_{i}.ID$ is the minimum from all the unsettled agents present at node $v$ then
6                         set $a_{i}.settled=1$
7                         update $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$
8                         if there is an agent $a_{j}$ from $G_{1}$ with $a_{j}.settled=0$ at node $v$ then
9                               update $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
10                  else
11                         set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
12                         if $a_{i}.prt\_out=a_{i}.prt\_in$ then
13                               set $a_{i}.state=backtrack$
14                        move through $a_{i}.prt\_out$
15            else if there is an agent $r_{j}$ at node $v$ with $r_{j}.settled=1$ then
16                   if $a_{i}.grp\_label=label$ and $a_{i}.dfs\_label=dfs\_label$, where $label$ and $dfs\_label$ are the component from $r_{j}^{v}(G_{2})$ then
17                         set $a_{i}.state=backtrack$
18                         set $a_{i}.prt\_out=a_{i}.prt\_in$
19                         move through $a_{i}.prt\_out$
20                  else if $a_{i}.grp\_label\neq label$ or $a_{i}.dfs\_label\neq dfs\_label$, where $label$ and $dfs\_label$ are the component from $r_{j}^{v}(G_{2})$ then
21                         set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
22                         if $a_{i}.prt\_out=a_{i}.prt\_in$ then
23                               set $a_{i}.state=backtrack$
24                               move through $a_{i}.prt\_out$
25                        else
26                               move through $a_{i}.prt\_out$
27      else if $a_{i}.success=False$ then
28             if  $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out=a_{i}.prt\_out$ then
29                   if $parent$ component of $r_{j}^{v}(G_{2})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
30                         if $a_{i}.prt\_out=\delta_{v}-1$ then
31                               $a_{i}.count_{2}=0$
32                               $a_{i}.skip=-1$
33                               $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
34                               $a_{i}.prt\_out=0$
35                               move through $a_{i}.prt\_out$
36                        else if $a_{i}.prt\_out\neq\delta_{v}-1$ then
37                               $a_{i}.count_{2}=0$
38                               $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{v}$
39                               move through $a_{i}.prt\_out$
40                  else if $parent$ component of $r_{j}^{v}(G_{2})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
41                         $a_{i}.count_{2}=0$
42                         $a_{i}.prt\_out=(a_{i}.prt\_out+1)$ mod $\delta_{v}$
43                         if $a_{i}.prt\_out=a_{i}.prt\_in$ then
44                               $a_{i}.state=backtrack$
45                               move through $a_{i}.prt\_out$
46                        else if $a_{i}.prt\_out\neq a_{i}.prt\_in$ then
47                               move through $a_{i}.prt\_out$
48            else if $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out\neq a_{i}.prt\_out$ or there is no agent from $G_{1}$ present at node $v$ then
49                   set $a_{i}.count_{2}=a_{i}.count_{2}+1$
50                   if $a_{i}.count_{2}<4n^{2}$ then
51                         move through $a_{i}.prt\_out$
52                  else if $a_{i}.count_{2}=4n^{2}$ then
53                         $a_{i}.count_{2}=0$ and $a_{i}.count_{3}=a_{i}.count_{3}+1$
54                         if $a_{i}.count_{3}<4n^{2}$ then
55                               set $a_{i}.state=explore$
56                               set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
57                               set $a_{i}.skip=a_{i}.prt\_out$
58                               set $a_{i}.prt\_out$ as the minimum port available at the current node except $a_{i}.skip$
59                               move through $a_{i}.prt\_out$
60                        else
61                               call $Group\_divide()$ (Algorithm 6)
62
63else if $r$ mod 2=0 then
64       if the movement of agent $a_{i}$ in round $r-1$ is successful then
65             set $a_{i}.success=True$
66      else
67             set $a_{i}.success=False$
Algorithm 4 $a_{i}.state=explore$
1 if $r$ mod 2 =1 then
2       if $a_{i}.success=True$ then
3             set $a_{i}.prt\_out=(a_{i}.prt\_in+1)\mod\delta_{v}$
4             set $a_{i}.count_{2}=0$
5             if  $parent$ component of $r_{j}^{v}(G_{2})$ is $-1$, where $r_{j}$ is the settled agent at node $v$ then
6                   if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
7                         $a_{i}.state=explore$
8                         $a_{i}.skip=-1$
9                         $a_{i}.prt\_out=0$
10                         $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
11                         move through $a_{i}.prt\_out$
12                  else if $a_{i}.prt\_out\neq$ minimum available port except $a_{i}.skip$ then
13                         if $a_{i}.prt\_out=a_{i}.skip$ then
14                               $a_{i}.prt\_out=(a_{i}.prt\_out+1)\mod\delta_{v}$
15                               if $a_{i}.prt\_out=$ minimum available port except $a_{i}.skip$ then
16                                     $a_{i}.state=explore$
17                                     $a_{i}.skip=-1$
18                                     $a_{i}.prt\_out=0$
19                                     $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
20                                     move through $a_{i}.prt\_out$
21                              else if $a_{i}.prt\_out\neq$ minimum available port except $a_{i}.skip$ then
22                                     $a_{i}.state=explore$
23                                     move through $a_{i}.prt\_out$
24                        else if $a_{i}.prt\_out\neq a_{i}.skip$ then
25                               $a_{i}.state=explore$
26                               move through $a_{i}.prt\_out$
27            else if $parent$ component of $r_{j}^{v}(G_{2})$ is not $-1$, where $r_{j}$ is the settled agent at node $v$ then
28                   if $a_{i}.prt\_out=parent$, where $parent$ is the component of $r_{j}^{v}(G_{2})$ then
29                         move through $a_{i}.prt\_out$
30                  else if $a_{i}.prt\_out\neq parent$, where $parent$ is the component of $r_{j}^{v}(G_{2})$ then
31                         $a_{i}.state=explore$
32                         move through $a_{i}.prt\_out$
33      else if $a_{i}.success=False$ then
34
35            if $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out=a_{i}.prt\_out$ then
36                   $a_{i}.count_{2}=0$
37                   set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
38                   set $a_{i}.state=explore$
39                   $a_{i}.skip=a_{i}.prt\_out$
40                   set $a_{i}.prt\_out$ as the minimum port available at the current node except $a_{i}.skip$
41                   move through $a_{i}.prt\_out$
42            else if $b_{i}$ from $G_{1}$ is present at node $v$ and $b_{i}.prt\_out\neq a_{i}.prt\_out$ or there is no agent from $G_{1}$ present at node $v$ then
43                   $a_{i}.count_{2}=a_{i}.count_{2}+1$
44                   if $a_{i}.count_{2}=4n^{2}$ then
45                         set $a_{i}.count_{2}=0$, and $a_{i}.count_{3}=a_{i}.count_{3}+1$
46                         if $a_{i}.count_{3}<4n^{2}$ then
47                               set $a_{i}.state=explore$
48                               set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
49                               set $a_{i}.skip=a_{i}.prt\_out$
50                               set $a_{i}.prt\_out$ as the minimum port available at the current node except $a_{i}.prt\_out$
51                               move through $a_{i}.prt\_out$
52                        else
53                               call $Group\_divide()$ (Algorithm 6)
54
55                  else if $a_{i}.count_{2}<4n^{2}$ then
56                         $a_{i}.count_{2}=a_{i}.count_{2}+1$
57                         move through $a_{i}.prt\_out$
58else if $r$ mod 2 = 0 then
59       if the movement of agent $a_{i}$ in round $r-1$ is successful then
60             set $a_{i}.success=True$
61      else
62             set $a_{i}.success=False$
Algorithm 5 $a_{i}.state=backtrack$
1 let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs
2 if $i\leq\lceil{\frac{x}{2}}\rceil$ then
[…]
4       $a_{i}.prt\_out=0$
5       set $a_{i}.state=explore$
6       $a_{i}.skip=-1$
7       set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
8       move through $a_{i}.prt\_out$
9else
[…]
11       $a_{i}.prt\_out=0$ set $a_{i}.state=explore$
12       $a_{i}.skip=-1$
13       set $a_{i}.dfs\_label=a_{i}.dfs\_label+1$
14       move through $a_{i}.prt\_out$
Algorithm 6 $Group\_divide()$
1 if $r$ mod 2=1 then
2       if unsettled agent $a_{j}$ from $G_{1}$ and $a_{j}.success=True$ then
3             if unsettled agents $a_{j}$ from $G_{1}$ is present at node $v$ with $a_{j}.state=backtrack$ then
4                   It does not do anything.
5            if unsettled agents $a_{j}$ from $G_{1}$ is present at node $v$ with $a_{j}.state=explore$ then
6                   if $label$ component of $a_{i}^{v}(G_{1})$ does not match with $a_{j}.grp\_label$ then
7                         set $a_{i}^{v}(G_{1}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
8                  else if $label$ component of $a_{i}^{v}(G_{1})$ matches with $a_{j}.grp\_label$ then
9                         It does not do anything
10      else if unsettled agent $a_{j}$ from $G_{1}$ and $a_{j}.success=False$ then
11             if unsettled agents $a_{j}$ from $G_{1}$ is present at node $v$ with $a_{j}.state=backtrack$ or $explore$ then
12                   if $a_{j}.count_{1}+1<16n^{2}$ then
13                         It does not do anything.
14                  else if $a_{j}.count_{1}+1=16n^{2}$ then
15                         let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs. Let $L=a_{i}.grp\_label=y$, $i\in[1,\,x]$
16                         if $i\leq\lceil{\frac{x}{2}}\rceil$ then
17                               set $a_{i}^{v}(G_{1}).(parent,label,dfs\_label)=(-1,\,L.0,\,a_{j}.dfs\_label+1)$
18                        else
19                               set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,L.1,\,a_{j}.dfs\_label+1)$
20
21else if $r$ mod 2=0 then
22       It does not do anything.
Algorithm 7 Algorithm for settled agent $a_{i}$ at node $v$ and at least one unsettled agent $a_{j}$ from $G_{1}$ present at node $v$
1 if $r$ mod 2=1 then
2       if unsettled agent $a_{j}$ from $G_{2}$ and $a_{j}.success=True$ then
3             if $a_{j}.state=backtrack$ then
4                   if $parent$ component of $a_{i}^{v}(G_{2})$ is $-1$ then
5                         if $(a_{j}.prt\_out+1)$ mod $\delta_{v}=$ minimum available port except for $a_{j}.skip$ then
6                               set $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,a_{j}.grp\_label,\,a_{j}.dfs\_label+1)$
7                        else if $(a_{j}.prt\_out+1)$ mod $\delta_{v}\neq$ minimum available port except for $a_{j}.skip$ then
8                               if $(a_{j}.prt\_out+1)$ mod $\delta_{v}=a_{j}.skip$ then
9                                     if $(a_{j}.prt\_out+2)=\delta_{v}-1$ then
10                                           set $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,a_{j}.grp\_label,\,a_{j}.dfs\_label+1)$
11                                    else if $(a_{j}.prt\_out+2)\neq\delta_{v}-1$ then
12                                           It does not do anything
13                              else if $(a_{j}.prt\_out+1)$ mod $\delta_{v}\neq a_{j}.skip$ then
14                                     It does not do anything
15                  else if $parent$ component of $a_{i}^{v}(G_{2})$ is not -1 then
16                         It does not do anything
17            if $a_{j}.state=explore$ then
18                   if $label$ and $dfs\_label$ components from $a_{i}^{v}(G_{2})$ do not match with $a_{j}.grp\_label$ and $a_{j}.dfs\_label$ then
19                         set $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(a_{j}.prt\_in,\,a_{j}.grp\_label,\,a_{j}.dfs\_label)$
20                  if $label$ and $dfs\_label$ components from $a_{i}^{v}(G_{2})$ match with $a_{j}.grp\_label$ and $a_{j}.dfs\_label$ then
21                         It does not do anything
22      else if unsettled agent $a_{j}$ from $G_{2}$ and $a_{j}.success=False$ then
23             if unsettled agents $b_{i}$ from $G_{1}$ is present at node $v$ then
24                   if $b_{i}.prt\_out=a_{j}.prt\_out$ and $a_{j}.state=backtrack$ then
25                         $a_{i}^{v}(G_{2}).(parent,\,label,\,dfs\_label)=(-1,\,a_{i}.grp\_label\,a_{i}.dfs\_label+1)$
26                  else if $b_{i}.prt\_out\neq a_{j}.prt\_out$ and $a_{j}.state=backtrack$ or there is no agent from $G_{1}$ then
27                         if $a_{j}.count_{2}+1=4n^{2}$ then
28                               if $a_{j}.count_{3}+1<4n^{2}$ then
29                                     set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,a_{j}.grp\_label,a_{j}.dfs\_label+1)$
30                              else if $a_{j}.count_{3}+1=4n^{2}$ then
31                                     let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs. Let $L=a_{j}.grp\_label$, $j\in[1,\,x]$
32                                     if $j\leq\lceil{\frac{x}{2}}\rceil$ then
33                                           set $a_{i}^{v}(G_{1}).(parent,label,dfs\_label)=(-1,\,L.0,\,a_{j}.dfs\_label+1)$
34                                    else
35                                           set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,L.1,\,a_{j}.dfs\_label+1)$
36                        else if $a_{j}.count_{2}+1<4n^{2}$ then
37                               It does not do anything
38                  else if $b_{i}.prt\_out=a_{j}.prt\_out$ and $a_{j}.state=explore$ then
39                         if $parent$ component of $a_{i}^{v}(G_{2})$ is $-1$ then
40                               if $a_{i}.prt\_out+1=\delta_{v}-1$ then
41                                     $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,a_{i}.grp\_label,\,a_{j}.dfs\_label+1)$
42                              else if $a_{i}.prt\_out+1\neq\delta_{v}-1$ then
43                                     It does not do anything
44                  else if $b_{i}.prt\_out\neq a_{j}.prt\_out$ and $a_{j}.state=explore$ or there is no agent from $G_{1}$ then
45                         if $a_{j}.count_{2}+1=4n^{2}$ then
46                               if $a_{j}.count_{3}+1<4n^{2}$ then
47                                     set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,a_{j}.grp\_label,a_{j}.dfs\_label+1)$
48                              else
49                                     let $\{a_{1},a_{2},...,a_{x}\}$ be the unsettled agents present at the current node in the increasing order of their IDs. Let $L=a_{j}.grp\_label$, $j\in[1,\,x]$
50                                     if $j\leq\lceil{\frac{x}{2}}\rceil$ then
51                                           set $a_{i}^{v}(G_{1}).(parent,label,dfs\_label)=(-1,\,L.0,\,a_{j}.dfs\_label+1)$
52                                    else
53                                           set $a_{i}^{v}(G_{2}).(parent,label,dfs\_label)=(-1,\,L.1,\,a_{j}.dfs\_label+1)$
54
55                        else if $a_{j}.count_{2}+1<4n^{2}$ then
56                               It does not do anything
57else if $r$ mod 2=0 then
58       It does not do anything.
Algorithm 8 Algorithm for settled agent $a_{i}$ at node $v$ and at least one unsettled agent $a_{j}$ from $G_{2}$ is present at node $v$